### v1

# An operator splitting analysis of Wasserstein–Fisher–Rao gradient flows

Abstract
Wasserstein-Fisher-Rao (WFR) gradient flows have been recently proposed as a powerful sampling tool that combines the advantages of pure Wasserstein (W) and pure Fisher-Rao (FR) gradient flows. Existing algorithmic developments implicitly make use of operator splitting techniques to numerically approximate the WFR partial differential equation, whereby the W flow is evaluated over a given step size and then the FR flow (or vice versa). This works investigates the impact of the order in which the W and FR operator are evaluated and aims to provide a quantitative analysis. Somewhat surprisingly, we show that with a judicious choice of step size and operator ordering, the split scheme can converge to the target distribution faster than the exact WFR flow (in terms of model time). We obtain variational formulae describing the evolution over one time step of both sequential splitting schemes and investigate in which settings the W-FR split should be preferred to the FR-W split. As a step towards this goal we show that the WFR gradient flow preserves log-concavity and obtain the first sharp decay bound for WFR.
1 Introduction
We consider the task of generating samples from a target probability distribution with density $\pi(x)\propto e^{-V_{\pi}(x)},\enskip x\in\mathbb{R}^{d}$ known up to a normalisation constant. Despite the conceptual simplicity of this task, its efficient implementation when $V_{\pi}$ is multi-modal, the underlying space is high dimensional and/or modes are separated by large distances remains challenging. Given the broad application of sampling to (Bayesian) statistics and statistical machine learning, several avenues have been investigated to derive efficient and scalable algorithms.
A natural way to formulate this task is via gradient flows [ambrosio2008gradient], which can be seen as optimisation of a functional measuring the dissimilarity to $\pi$, typically the Kullback–Leibler (KL) divergence [wibisono2018sampling]. This formulation yields considerable freedom in the design of sampling algorithms; arguably the most well-known being the so-called Wasserstein Gradient flows, hereafter W flows. It is well-known that when $\pi$ satisfies a Log-Sobolev Inequality (LSI), the W flow converges exponentially fast to $\pi$, with rate depending on the Log-Sobolev constant. This highlights an inherent limitation of W flows, namely that when the LSI constant is large (e.g. as is typically the case in multi-modal densities with well separated modes), the convergence rate can be prohibitively slow [schlichting2019poincare].
Recent research efforts have instead considered gradient flows in the Fisher-Rao geometry (FR flow). These are well known in the biological literature as describing the macroscopic properties of a population with varying traits or species, and is sometimes referred to as birth-death or replicator dynamics [kimura_stochastic_1965, schuster1983replicator, Cressman2006]. It is now known that FR flows are intimately connected to mirror descent [chopin2023connection], stochastic filtering [akyildiz2017probabilistic, halder_gradient_2017, Pathiraja2024, DelMoral1997] and Sequential Monte Carlo [us]. The have also recently been exploited to develop sampling algorithms [Nusken2024, maurais2024sampling, Wang2024, chen2023sampling, lu2023birth, Lu2019], most notably due to the fact that it is possible to achieve convergence rates independent of the properties of $V_{\pi}$ [carrillo_fisher-rao_2024, Lu2019]. However, the development of stable and efficient numerical approximations of pure FR flows remains a challenge.
In this work, we consider the so-called Wasserstein–Fisher–Rao (WFR) gradient flow, wherein the metric is given by the direct sum of the Wasserstein and Fisher-Rao metrics. We refer the reader to [liero_optimal_2018] for a rigorous treatment of this metric and its corresponding gradient flow. WFR gradient flows combine the diffusive behaviour of W flows with the birth-death or reactive properties of FR flows and enjoy better convergence properties than both W and FR alone [Lu2019, chen2023sampling]. They have a natural interpretation as combining ‘exploration’ or ‘mutation’ to provide new particles (W flow) with ‘selection’ where particles that are a poor fit to the target are killed [Pathiraja2024]. The convergence properties of WFR flows have been theorised to improve on both W and FR flows, yet the best known results are obtained under strong conditions on the ration between $\pi$ and the initial distribution $\mu_{0}$ [Lu2019, Appendix B] and [lu2023birth, Remark 2.6] and, as we show later, are not sharp. Research into WFR flows is active; several open questions and areas of investigation remain both in terms of algorithmic development and theoretical investigation.
The structure of WFR gradient flows naturally lends itself to numerical methods built on operator splitting [Hundsdorfer2003]. Splitting of the WFR PDE has been considered in several papers when deriving algorithms [us, lu2023birth, Lu2019] but without a precise characterisation of the effect of the splitting choice.
Operator splitting has been previously exploited to build splitting schemes for the kinetic Langevin SDE arising from the W flow to obtain algorithms with no time discretisation bias [chada2023unbiased] and methods with higher accuracy for fixed computational cost [leimkuhler2013rational, bou2010long]. Our work aims to show that splitting schemes can exhibit faster convergence to the target than the exact dynamics and thus a wise choice of splitting can reduce computational cost.
In this work we focus on exact splitting, i.e. we do not consider numerical schemes to approximately solve the W and FR flow and their numerical errors but rather focus on the numerical error introduced by time discretisations via operator splitting. We demonstrate that numerical error introduced by operator splitting can be exploited to accelerate convergence to equilibrium. A particular advantage of this observation is that ordering of the operators does not affect computational cost, and thus speed up can be obtained at no additional cost and without introducing additional numerical approximations.
This is particularly relevant when the computational budget is limited, as the speed up due to splitting is mostly observed in the early iterations. We thus claim that the continuous time process need not be the target of our numerical approximations but rather the splitting scheme which gives us the fastest convergence given $\pi,\mu_{0}$. This then paves the way for the development of practical numerical algorithms (i.e. space-time discretisations) that approximate the operator split WFR flow rather than the fully coupled WFR gradient flow. There are a myriad schemes for numerically approximating both W [NEURIPS2019_6a8018b3_salim, salim2020wasserstein, mou2021high] and FR [nusken2024transport, maurais2024sampling, Wang2024, korba2022adaptive, chen2023sampling] flows and we leave their detailed investigation to future work.
Our main contributions are as follows:
  • We derive variational formulae (PDEs) that quantify the impact of operator splitting as function of the step size (Section 2) and demonstrate that a carefully chosen operator splitting method can yield theoretical convergence rates that are faster than the fully continuous time WFR flow (Section 6) without increasing the computational cost.
  • To do so, we obtain the first convergence result for the continuous time WFR flow which clearly shows that the rate of convergence is the sum of the rate of the W flow and that of the FR flow (Section 5). This result is obtained after establishing that the WFR preserves log-concavity thanks to the strong regularising properties of the FR flow (Section 4).
  • We provide as motivation a precise quantitative characterisation in the multivariate Gaussian case (Section 3) which allows us to gain some intuition on general guidelines on the choice of initial distribution $\mu_{0}$ to obtain faster convergence (Section 6).
Notation
We define some notation that will be used throughout the manuscript. For all differentiable functions $f$ we denote the gradient by $\nabla f$. Furthermore, if $f$ is twice differentiable we denote by $\nabla^{2}f$ its Hessian and by $\Delta f$ its Laplacian. We denote by $\mathcal{P}(^{d})$ the set of probability measures over $\mathcal{B}(^{d})$, and endow this space with the topology of weak convergence. We denote by $\mathcal{P}_{2}^{ac}(^{d})$ the manifold of absolutely continuous probability measures on d with finite second moment. Every $p\in\mathcal{P}_{2}^{ac}(^{d})$ will be identified with its (Lebesgue) density $p(x)$. At a $p\in\mathcal{P}_{2}^{ac}(^{d})$ such that $p(x)>0$ everywhere, the tangent space of $\mathcal{P}_{2}^{ac}(^{d})$ is given by $\mathcal{T}=\{\sigma\in C^{\infty}(\mathbb{R}^{d}):\int\sigma(x)\,\mathrm{d}x=0\}$. If $p$ has support $\mathrm{supp}(p)\subsetneq\mathbb{R}^{d}$, then the tangent space is $\mathcal{T}=\{\sigma\in C^{\infty}(\mathbb{R}^{d}):\sigma|_{\mathrm{supp}(p)^{C}}=0,\int\sigma(x)\,\mathrm{d}x=0\}$. Throughout this manuscript we denote the target by $\pi$ and assume $\pi(x)\propto e^{-V_{\pi}(x)}$. The initial distribution of each flow is denoted by $\mu_{0}$. The Kullback–Leibler divergence is defined for $\nu,\mu$ admitting a density w.r.t. Lebesgue as $\mathrm{KL}(\nu||\mu)=\int\log(\nicefrac{{\nu(x)}}{{\mu(x)}})\nu(x)dx$.
2 Operator Splitting for the WFR PDE
The gradient flow of $\mathrm{KL}(\mu||\pi)$ w.r.t. the geometry induced by the Wasserstein–Fisher–Rao distance is given by the following PDE [Lu2019, Theorem 3.1]
$\partial_{t}\mu_{t}=f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}),$ (2.1)
where
$f_{\textrm{W}}(\mu)$ $:=\nabla\cdot(\mu\nabla g(\mu)),$ (2.2)
$f_{\textrm{FR}}(\mu)$ $:=-\mu\left(g(\mu)-\mathbb{E}_{\mu}\left[g(\mu)\right]\right),$ (2.3)
$g(\mu)$ $:=\log\frac{\mu}{\pi}.$ (2.4)
and $f_{\textrm{W}}$ and $f_{\textrm{FR}}$ correspond to the Wasserstein and Fisher–Rao operators respectively, and takes the form of a reaction-diffusion equation. This PDE is a combination of the Wasserstein (W) flow and of the Fisher–Rao (FR) flow and enjoys better convergence properties than both W and FR alone [Lu2019, chen2023sampling].
The WFR PDE (2.1) corresponds to the underlying continuous time dynamics of many recent sampling methods as highlighted in [us]. Recent methods involve using a variety of numerical schemes e.g. interacting particle based methods [Lu2019, lu2023birth, chen2023sampling] and kernel based methods [maurais2024sampling, Nusken2024, Wang2024]. A well-known approach to discretising PDEs involving sums of operators is operator splitting [Hundsdorfer2003, blanes2024splitting] which has been applied to the W flow to split advection and diffusion and led to a vast number of numerical approximations of Langevin dynamics [abdulle2015long, duncan2017nonreversible, chada2023unbiased, leimkuhler2013rational, bou2010long]. Such operator splitting is particularly natural for the WFR PDE, where the W flow is characterised by advection and diffusion operators whilst the FR flow is characterised by reaction dynamics. Splitting of the WFR PDE has been considered in several papers when deriving algorithms [us, lu2023birth, Lu2019] but without a precise characterisation of the effect of the splitting choice.
Throughout this manuscript, we focus on time discretisation of (2.1) via operator splitting under the assumption that the operators are evaluated exactly over a single time step, i.e. without further discretisations using either determinstic or Monte Carlo type approximations. Our motivation for doing so is to quantify biases due solely to splitting error, thereby giving rise to a new PDE which can potentially yield faster (continuous time) convergence rates than the original WFR PDE. This opens the door for the development and study of numerical schemes whose underlying dynamics correspond to the split WFR flow rather than its fully continuous time counterpart. In particular, we assume that the solution operator of the W flow for a time step $\gamma$, $S_{\textrm{W}}(\gamma,v)(x)$, can be evaluated exactly. Similarly, we assume the solution operator of the FR flow given by (e.g. [chen2023sampling, Appendix B.1])
$S_{\textrm{FR}}(\gamma,v)(x)\propto\pi(x)^{1-e^{-\gamma}}v(x)^{e^{-\gamma}},$ (2.5)
can also be evaluated exactly. A commonly used first order accurate splitting method is sequential splitting or Lie-Trotter splitting (see, e.g., [blanes2024splitting] for a recent review), where each operator is solved over a step size in an alternating manner. More specifically, consider a time-discretisation $t_{0}=0<t_{1}<t_{2}\cdots<t_{M}=T$ with $t_{n+1}-t_{n}=\gamma\enskip\forall\enskip i=1,2,\dots M$. A sequential splitting applied to (2.1) takes the form
$\hat{\nu}_{1}(x;\gamma)$ $=S_{\textrm{W}}(\gamma,\mu_{0})(x)$ (2.6)
$\nu_{1}(x;\gamma)$ $=S_{\textrm{FR}}(\gamma,\hat{\nu}_{1})(x)$
$\hat{\nu}_{2}(x;\gamma)$ $=S_{\textrm{W}}(\gamma,\nu_{1})(x)$
$\nu_{2}(x;\gamma)$ $=S_{\textrm{FR}}(\gamma,\hat{\nu}_{2})(x)$
$\vdots$
where $S_{\textrm{FR}}(\gamma,v)$ and $S_{\textrm{W}}(\gamma,v)$ denote the solution operators corresponding to $f_{\textrm{FR}}(\mu)$ and $f_{\textrm{W}}(\mu)$ acting on $v$ over time interval of size $\gamma$, respectively. We refer to (2.6) as Wasserstein then Fisher–Rao, or W-FR for short. It is known that the $n$th iterate of the splitting scheme (2.6), $\nu_{n}$, converges to the solution of (2.1) at time $t=n\gamma$, $\mu_{n\gamma}$, as $\gamma\rightarrow 0$ (at rate $\mathcal{O}(\gamma)$) under mild conditions on $f_{\textrm{FR}}$, $f_{\textrm{W}}$ and $\pi$ (see e.g. [Hundsdorfer2003]).
The splitting scheme in (2.6) corresponds to evaluating the W solution operator first. A valid alternative sequential splitting scheme to (2.6) involves instead evaluating the FR solution operator first, i.e.
$\hat{\eta}_{1}(x;\gamma)$ $=S_{\textrm{FR}}(\gamma,\mu_{0})(x)$ (2.7)
$\eta_{1}(x;\gamma)$ $=S_{\textrm{W}}(\gamma,\hat{\eta}_{1})(x)$
$\hat{\eta}_{2}(x;\gamma)$ $=S_{\textrm{FR}}(\gamma,\eta_{1})(x)$
$\eta_{2}(x;\gamma)$ $=S_{\textrm{W}}(\gamma,\hat{\eta}_{2})(x)$
$\vdots$
We refer to (2.7) as Fisher–Rao then Wasserstein, or FR-W for short. When the operators $S_{\textrm{FR}}$ and $S_{\textrm{W}}$ do not commute, it is known that $\eta_{i}(x;\gamma)\neq\nu_{i}(x;\gamma)$ for $\gamma>0$ and $i=1,2,3,\dots$ [Hundsdorfer2003]. It is expected that for most target distributions, $S_{\textrm{FR}}$ and $S_{\textrm{W}}$ will not commute; the following example demonstrates this occurs even in the multivariate Gaussian case.
In the remainder of the manuscript, we rely on the following assumptions on the target distribution $\pi$ and the distribution at which the WFR PDE (2.1) is initialised $\mu_{0}$ in developing analytic results.
Assumption 1.
The following conditions on the initial and target distribution hold
  1. (a)
    $\pi(x)\propto e^{-V_{\pi}(x)}$ where $V_{\pi}$ is continuously differentiable with bounded Hessian, and satisfies a Log-Sobolev inequality (LSI): for all $\mu\in\mathcal{P}_{2}^{ac}(^{d})$
    $\mathrm{KL}(\mu||\pi)\leq\frac{\lambda_{\pi}}{2}\int\mu\left\lvert\nabla\log\frac{\mu}{\pi}\right\rvert^{2}:=\frac{\lambda_{\pi}}{2}\mathcal{I}(\mu||\pi),$ (2.8)
    where $\lambda_{\pi}<\infty$ denotes the Log-Sobolev constant of $\pi$ and is $\mathcal{I}(\mu||\pi)$ the relative Fisher information between $\mu$ and $\pi$;
  2. (b)
    $\mu_{0}(x)\propto e^{-V_{0}(x)}$ where $V_{0}$ is continuously differentiable with bounded derivative, satisfies an LSI with constant $\lambda_{0}<\infty$ and $\mu_{0}\ll\pi$.
Assumption 1(a) implies that the Wasserstein flow targeting $\pi$, $\partial_{t}\mu_{t}=f_{\textrm{W}}(\mu_{t})$, is well defined and the solution of the corresponding PDE is absolutely continuous w.r.t. $\pi$ for all $t\geq 0$. The bounded Hessian requirement implies that $\nabla V_{\pi}$ is Lipschitz continuous, we denote the Lipschitz constant by $L_{\pi}$. Assumption 1(b) is mild since $\mu_{0}$ is often user-chosen and guarantees that $\nabla V_{0}$ is Lipschitz continuous with constant $L_{0}$.
We further assume that the FR PDE $\partial_{t}\mu_{t}=f_{\textrm{FR}}(\mu_{t})$ and WFR PDE $\partial_{t}\mu_{t}=f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t})$ are well-posed and that their solutions, where they exist, are absolutely continuous w.r.t. $\pi$. These assumptions guarantee that the solution operators $S_{\textrm{W}},S_{\textrm{FR}}$ are well-defined.
Remark 2.1.
One could also apply a splitting scheme for pure W to split the advection and diffusion parts (leading to proximal samplers, see e.g. [wibisono2018sampling, Section 4.2]). Similarly, one could also apply a splitting scheme to the pure FR flow by identifying an exploration step and an exploitation step as in [chen_efficient_2024, Section 4]. We leave analysis of these alternative splitting schemes for future work.
2.1 Variational Formula: W-FR
We begin by analysing the evolution obtained by using the splitting scheme in (2.6). In particular, we obtain a PDE which describes the evolution of the splitting solution over time interval $0,\gamma]$. The advantage of this characterisation is that we can identify a perturbation to the exact dynamics and more easily characterise how splitting biases modify convergence speed. The traditional route to analysing decay of $\mathrm{KL}$ of the WFR dynamics in e.g. [us, Proposition 3.1], [lu2023birth, Remark 2.6], [Lu2019, Appendix B] is not suited to the splitting schemes as it does not allow us to distinguish the order in which the W and FR solution operators are applied.
The following proposition quantifies the variation of $\nu_{\gamma}$ with respect to the step size $\gamma$ using standard variational calculus tools. For the sake of notational brevity, we rely on the shorthand notation $\nu_{\gamma}$ to refer to $\nu_{1}(x;\gamma)$.
Proposition 2.2.
Sequential split (W-FR) PDE: Let us denote $\nu_{\gamma}=\nu_{1}(x;\gamma)$ and let Assumption 1 hold. The variation of one sequential split step in the order W-FR (2.6) of size $\gamma$ corresponds to the PDE
$\partial_{\gamma}\nu_{\gamma}=f_{\textrm{W}}(\nu_{\gamma})+f_{\textrm{FR}}(\nu_{\gamma})+(e^{\gamma}-1)\nu_{\gamma}(|\nabla g(\nu_{\gamma})|^{2}-\mathbb{E}_{\nu_{\gamma}}[|\nabla g|^{2}]),$ (2.9)
where $g$ is as defined in (2.4).
Proof.
Here we use the notation $\rho_{\gamma}$ to refer to the unnormalised form of $\nu_{\gamma}$. Denote by $f_{\textrm{G}}$ the unnormalised FR operator,
$f_{\textrm{G}}(\mu):=\mu\log\frac{\pi}{\mu}.$
The unnormalised sequential split solution (in the order W-FR) at time $\gamma$ is given by
$\rho_{\gamma}=S_{\textrm{G}}(\gamma,S_{\textrm{W}}(\gamma,\mu_{0})),$
where $S_{\textrm{G}}(\gamma,v)$ denotes the abstract solution operator of $\partial_{\gamma}\rho_{\gamma}=f_{\textrm{G}}(\rho_{\gamma})$ initialised at $v$ after time $\gamma$ and $S_{\textrm{W}}(\gamma,v)$ denotes the solution operator of $\partial_{\gamma}\rho_{\gamma}=f_{\textrm{W}}(\rho_{\gamma})$. Using the chain rule, the first variation of $\rho_{\gamma}$ with respect to $\gamma$ is (see also (3.4) in [descombes_lietrotter_2013])
$\partial_{\gamma}\rho_{\gamma}=f_{\textrm{G}}(\rho_{\gamma})+\partial_{2}S_{\textrm{G}}(\gamma,v_{\gamma})f_{\textrm{W}}(v_{\gamma}),$ (2.10)
where $v_{\gamma}:=S_{\textrm{W}}(\gamma,\mu_{0})$ and $\partial_{i}$ denotes the directional derivative w.r.t. the $i$-th argument. Due to [chen2023sampling, Appendix B.1], the solution operator $S_{\textrm{G}}$ has an explicit form given by $S_{\textrm{G}}(\gamma,v)=\pi^{1-e^{-\gamma}}v^{e^{-\gamma}}$, so that $\partial_{2}S_{\textrm{G}}(\gamma,v)h$, i.e. the directional derivative w.r.t. the second argument $v\in\mathcal{P}_{2}^{ac}$ in the direction $h\in\mathcal{T}$ is easily computed as,
$\partial_{2}S_{\textrm{G}}(\gamma,v)h=\left.\frac{d}{d\epsilon}S_{\textrm{G}}(\gamma,v+\epsilon h)\right|_{\epsilon=0}$ $=\left.\frac{d}{d\epsilon}\pi^{1-e^{-\gamma}}(v+\epsilon h)^{e^{-\gamma}}\right|_{\epsilon=0}$
$=e^{-\gamma}\pi^{1-e^{-\gamma}}v^{e^{-\gamma}-1}h$
$=e^{-t}S_{\textrm{G}}(t,v)hv^{-1}.$
It is easy to see that $\partial_{2}S_{\textrm{G}}(\gamma,v)h$ is a linear operator and that $x\mapsto\partial_{2}S_{\textrm{G}}(\gamma,v)(x)=e^{-t}(\pi(x)/v(x))^{1-e^{-t}}$ is continuous if $\pi,v$ are continuous and $v(x)>0$. Substituting back into (2.10) yields
$\partial_{\gamma}\rho_{\gamma}=f_{\textrm{G}}(\rho_{\gamma})+e^{-\gamma}S_{\textrm{G}}(\gamma,v_{\gamma})f_{\textrm{W}}(v_{\gamma})v_{\gamma}^{-1}.$ (2.11)
Notice that the r.h.s can be written entirely in terms of $\rho_{\gamma}$ since $S_{\textrm{G}}$ is an invertible operator, with inverse given by
$S_{\textrm{G}}^{-1}(\gamma,v)=v=\exp(e^{\gamma}\log(\pi^{-1+e^{-\gamma}}S_{\textrm{G}}(\gamma,v)))=\pi^{1-e^{\gamma}}S_{\textrm{G}}(\gamma,v)^{e^{\gamma}},$
Since $\rho_{\gamma}=S_{\textrm{G}}(\gamma,v_{\gamma})$ where $v_{\gamma}=S_{\textrm{W}}(\gamma,\mu_{0})$, we have using the above that,
$v_{\gamma}=\pi^{1-e^{\gamma}}\rho_{\gamma}^{e^{\gamma}}.$
Then $f_{\textrm{W}}(v_{\gamma})$ can be expressed in terms of $\rho_{\gamma}$ since, using the shorthand notation $w:=\nabla\log\frac{\rho_{\gamma}}{\pi}$,
$f_{\textrm{W}}(v_{\gamma})$ $=\nabla\cdot\left(v_{\gamma}\nabla\log\frac{v_{\gamma}}{\pi}\right)$
$=e^{\gamma}\nabla\cdot\left(\pi^{-e^{\gamma}+1}\rho_{\gamma}^{e^{\gamma}}w\right)$
$=e^{\gamma}(\pi^{-e^{\gamma}+1}\nabla\cdot\left(\rho_{\gamma}^{e^{\gamma}-1}\rho_{\gamma}w\right)+\rho_{\gamma}^{e^{\gamma}}w\cdot\nabla(\pi^{-e^{\gamma}+1}))$
$=e^{\gamma}(\pi^{-e^{\gamma}+1}\rho_{\gamma}^{e^{\gamma}-1}\nabla\cdot\left(\rho_{\gamma}w\right)+\pi^{-e^{\gamma}+1}\rho_{\gamma}w\cdot\nabla(\rho_{\gamma}^{e^{\gamma}-1})+\rho_{\gamma}^{e^{\gamma}}w\cdot\nabla(\pi^{-e^{\gamma}+1}))$
$=e^{\gamma}\pi^{-e^{\gamma}+1}\rho_{\gamma}^{e^{\gamma}-1}(f_{\textrm{W}}(\rho_{\gamma})+\underbrace{\rho_{\gamma}^{2-e^{\gamma}}w\cdot\nabla(\rho_{\gamma}^{e^{\gamma}-1})+\rho_{\gamma}\pi^{e^{\gamma}-1}w\cdot\nabla(\pi^{-e^{\gamma}+1})}_{=:R}),$
and notice that $R$ can be simplified further as
$R$ $=\rho_{\gamma}\left(\rho_{\gamma}^{1-e^{\gamma}}w\cdot\nabla(\rho_{\gamma}^{e^{\gamma}-1})+\pi^{e^{\gamma}-1}w\cdot\nabla(\pi^{-e^{\gamma}+1})\right)$
$=\rho_{\gamma}\left((e^{\gamma}-1)\rho_{\gamma}^{-1}w\cdot\nabla\rho_{\gamma}-w\cdot(e^{\gamma}-1)\pi^{-1}\nabla\pi+\nabla\cdot w\right)$
$=(e^{\gamma}-1)\rho_{\gamma}|w|^{2},$
so that
$f_{\textrm{W}}(v_{\gamma})=e^{\gamma}\pi^{-e^{\gamma}+1}\rho_{\gamma}^{e^{\gamma}-1}(f_{\textrm{W}}(\rho_{\gamma})+(e^{\gamma}-1)\rho_{\gamma}|w|^{2}).$
Substituting all the derived expressions back into (2.11) yields
$\partial_{\gamma}\rho_{\gamma}$ $=f_{\textrm{G}}(\rho_{\gamma})+e^{-\gamma}e^{\gamma}\pi^{-e^{\gamma}+1}\rho_{\gamma}^{e^{\gamma}-1}(f_{\textrm{W}}(\rho_{\gamma})+(e^{\gamma}-1)\rho_{\gamma}|w|^{2})\pi^{e^{\gamma}-1}\cdot\rho_{\gamma}^{1-e^{\gamma}}$
$=f_{\textrm{W}}(\rho_{\gamma})+f_{\textrm{G}}(\rho_{\gamma})+(e^{\gamma}-1)\rho_{\gamma}|w|^{2}.$
Then by an application of [us, Lemma A.1] to obtain the PDE for the normalised density $\nu_{\gamma}$, the claim holds.
The above proposition shows that splitting introduces a perturbation to the WFR PDE (2.1) of the form
$f_{\textrm{P}}(\nu_{\gamma}):=\nu_{\gamma}\left(\left|\nabla g(\nu_{\gamma})\right|^{2}-\mathbb{E}_{\nu_{\gamma}}\left[\left|\nabla g(\nu_{\gamma})\right|^{2}\right]\right),$ (2.12)
where $g$ is as defined in (2.4), whose contribution is modulated by an additional factor $(e^{\gamma}-1)$ which increases monotonically with step size $\gamma$. Somewhat remarkably, a simple splitting procedure is sufficient to generate an acceleration under certain conditions as we will show in Sections 3 and 6.
2.2 Variational Formula: FR-W
The reverse scheme described in (2.7) requires a more involved analysis than in Proposition 2.2, due to the lack of invertibility of $S_{\textrm{W}}$ in general. We instead rely on techniques used in developing Alekseev-Groebner type formulae, in particular, the formal Lie calculus, as is common in the analysis of operator splitting schemes [descombes_lietrotter_2013, Hundsdorfer2003]. Once again, we rely on the shorthand notation $\eta_{\gamma}$ to refer to $\eta_{1}(x;\gamma)$.
Proposition 2.3.
Sequential split PDE: FR-W. Let us denote $\eta_{\gamma}=\eta_{1}(x;\gamma)$ and let Assumption 1 hold. The variation of one sequential split step of size $\gamma$ in the order FR-W (2.7) corresponds to the PDE
$\partial_{\gamma}\eta_{\gamma}=f_{\textrm{W}}(\eta_{\gamma})+f_{\textrm{FR}}(\eta_{\gamma})+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,-f_{\textrm{P}}(\eta_{\tau})\right)d\tau,$ (2.13)
where $f_{\textrm{P}}$ is as defined in (2.12). An alternative representation is given by
$\partial_{\gamma}\eta_{\gamma}=f_{\textrm{W}}(\eta_{\gamma})+f_{\textrm{FR}}(\eta_{\gamma})+\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\eta_{\gamma}\left(g_{k}(\eta_{\gamma})-\mathbb{E}_{\eta_{\gamma}}[g_{k}]\right),$ (2.14)
where
$g_{0}(\eta_{\gamma})$ $=-g(\eta_{\gamma})=-\log\frac{\eta_{\gamma}}{\pi},$
$g_{k}(\eta_{\gamma})$ $=-g_{k-1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})+\nabla\log\frac{\eta_{\gamma}}{\pi}\cdot\nabla g_{k-1}(\eta_{\gamma})+\frac{1}{\eta_{\gamma}}\nabla\cdot(\eta_{\gamma}\nabla g_{k-1}(\eta_{\gamma})),\quad\text{for}\enskip k=2,3,\dots$
and $g_{k-1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})$ denotes the directional derivative of $g_{k-1}$ at $\eta_{\gamma}$ in the direction $f_{\textrm{W}}(\eta_{\gamma})$.
Proof.
By similar reasoning as in Section 2.1, the sequential split solution at time $\gamma$ in the order FR-W is given by $\eta_{\gamma}=S_{\textrm{W}}(\gamma,S_{\textrm{FR}}(\gamma,\mu_{0}))$ and its first variation w.r.t. $\gamma$ can be computed as
$\partial_{\gamma}\eta_{\gamma}=f_{\textrm{W}}(\eta_{\gamma})+\partial_{2}S_{\textrm{W}}(\gamma,u_{\gamma})f_{\textrm{FR}}(u_{\gamma}),$ (2.15)
where $u_{\gamma}:=S_{\textrm{FR}}(\gamma,\mu_{0})$. By Lemma 1 in [descombes_lietrotter_2013] and recalling that $S_{\textrm{W}}$ is a linear operator for which $\partial_{2}S_{\textrm{W}}(t,v)h=S_{\textrm{W}}(t,h)$, we have
$\partial_{2}S_{\textrm{W}}(\gamma,u_{\gamma})f_{\textrm{FR}}(u_{\gamma})$ $=f_{\textrm{FR}}(S_{\textrm{W}}(\gamma,u_{\gamma}))+\int_{0}^{\gamma}\partial_{2}S_{\textrm{W}}(\gamma-\tau,u_{\gamma})[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,u_{\gamma}))d\tau$
$=f_{\textrm{FR}}(\eta_{\gamma})+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,u_{\gamma}))\right)d\tau,$ (2.16)
since $S_{\textrm{W}}$ is a linear operator and where $f_{\textrm{W}},f_{\textrm{FR}}](v)$ refers to the Lie commutator (see (3.2a) in [descombes_lietrotter_2013]),
$[f_{\textrm{W}},f_{\textrm{FR}}](v)$ $=D_{\textrm{FR}}f_{\textrm{W}}(v)-D_{\textrm{W}}f_{\textrm{FR}}(v),$ (2.17)
$=f_{\textrm{W}}^{\prime}f_{\textrm{FR}}(v)-f_{\textrm{FR}}^{\prime}f_{\textrm{W}}(v)=v\left(g_{1}(v)-\mathbb{E}_{v}\left[g_{1}\right]\right),$ (2.18)
where $D_{\textrm{FR}}$ is the Lie derivative operator associated with $f_{\textrm{FR}}$, i.e. $D_{\textrm{FR}}H(v)=H^{\prime}(v)f_{\textrm{FR}}(v)$ for any unbounded nonlinear operator $H:X\rightarrow\mathcal{T}$, $X\subset\mathcal{P}_{2}^{ac}$ and $H^{\prime}(v)$ denotes the Fréchet derivative operator evaluated at $v$. We note that while $f_{\textrm{W}}$ and $f_{\textrm{FR}}$ may not be Fréchet differentiable at any arbitrary $v$, their directional derivative exists and is sufficient to define the Lie derivative, therefore we work with the directional derivative. With a slight abuse of notation, we use $f_{\textrm{W}}^{\prime}f_{\textrm{FR}}(v)$ in place of $f_{\textrm{W}}^{\prime}(v)f_{\textrm{FR}}(v)$. Substituting (2.16) into (2.15) yields that the variation of $\eta_{\gamma}$ takes the form of the usual WFR PDE plus a perturbation term,
$\partial_{\gamma}\eta_{\gamma}=f_{\textrm{W}}(\eta_{\gamma})+f_{\textrm{FR}}(\eta_{\gamma})+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,u_{\gamma}))\right)d\tau,$ (2.19)
where $f_{\textrm{W}},f_{\textrm{FR}}]S_{\textrm{W}}(\tau,u_{\gamma})\in\mathcal{T}$. Substituting in (A.6) developed in Lemma A.2 for the Lie commutator leads to (2.13).
Finally, the perturbation term in (2.19) is characterised in Lemma A.1 using the repeated commutators $\mathcal{G}^{k}$ defined therein,
$\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,u_{\gamma}))\right)d\tau=\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}^{k}(\eta_{\gamma}).$
The result then follows using Lemma A.2. ∎
The expression in (2.14) is obtained by writing the perturbation term in (2.13) using the Lie commutator in (2.17) and higher order commutators. We work with this formulation as the inverse of the solution operator does not lead to a tractable form, when it exists. A similar approach can be used to obtain the perturbation term in Proposition 2.2 with $g_{k}=|\nabla g|^{2}$ for all $k=1,2,3,\dots$. We do not use this approach in the proof of Proposition 2.2 but exploit the invertibility of the unnormalised FR solution operator.
As Proposition 2.3 shows, the evolution of the splitting solution obtained with FR-W does not correspond to that obtained in Proposition 2.2 for W-FR. In the multivariate Gaussian setting, i.e. with $\mu_{0}(x)=\mathcal{N}(x;m_{0},C_{0})$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$, a more interpretable PDE reminiscent of (2.9) can be obtained (see Appendix B.2 for a derivation),
$\partial_{\gamma}\eta_{\gamma}$ $=f_{\textrm{W}}(\eta_{\gamma})+f_{\textrm{FR}}(\eta_{\gamma})-\eta_{\gamma}\left(g_{\text{FRW}}(\eta_{\gamma})-\mathbb{E}_{\eta_{\gamma}}[g_{\text{FRW}}(\eta)]\right),$ (2.20)
$g_{\text{FRW}}(\eta)$ $:=\nabla g(\eta)^{\top}\frac{C_{\pi}}{2}\left(e^{2\gamma C_{\pi}^{-1}}-I\right)\nabla g(\eta).$
We investigate the multivariate Gaussian case in further detail in the next section.
3 Splitting & convergence speed: multivariate Gaussian case
The multivariate Gaussian case offers a tractable setting in which to obtain quantitative results on the impact of operator order in splitting of the WFR flow. More specifically, consider the case $\mu_{0}(x)=\mathcal{N}(x;m_{0},C_{0})$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$ with $C_{\pi}\neq C_{0}$. It can be readily shown that the solution of the WFR PDE (2.1) at any $t>0$ is also Gaussian $\mu_{t}(x)=\mathcal{N}(x;m_{t},C_{t})$ with moments given by
$C_{t}$ $=C_{\pi}+(e^{\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}]e^{\Gamma t}-(2I+C_{\pi})^{-1})^{-1},$ (3.1)
$m_{t}$ $=m_{\pi}+(C_{t}-C_{\pi})e^{tC_{\pi}^{-1}}(C_{0}-C_{\pi})^{-1}(m_{0}-m_{\pi}),$ (3.2)
where $\Gamma:=C_{\pi}^{-1}+\frac{1}{2}I$. A derivation of (3.1)-(3.2) can be found in Lemma B.2 (letting $M_{t}=0$), which extends the derivation of $C_{t}^{-1}$ provided in [liero2025evolution]. To the best of our knowledge, the analytic solution for $m_{t}$ provided in (3.2) has not yet appeared in the literature.
Similarly, the splitting schemes yield Gaussian solutions at each time, as it is well-known that both W and FR flows preserve Gaussianity. We denote by $b_{n\gamma},Q_{n\gamma}$ the mean and covariance respectively at step $n$ of the W-FR split scheme with step size $\gamma$ and $a_{n\gamma},P_{n\gamma}$ the mean and covariance at step $n$ of the FR-W split. We write $b_{\gamma},Q_{\gamma}$ and $a_{\gamma},P_{\gamma}$ when $n=1$.
The similarity in structure of the sequential split PDEs (2.9) and (2.20) in the Gaussian setting also allows for obtaining analytic solutions of the mean and covariance in a similar fashion. By an application of Lemma B.2 with $\gamma$ in place of $t$ and setting $M_{\gamma}=(e^{\gamma}-1)I$, one obtains
$Q_{\gamma}$ $=C_{\pi}+(e^{\Gamma\gamma}(C_{0}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1})^{-1},$ (3.3)
$b_{\gamma}$ $=m_{\pi}+(Q_{\gamma}-C_{\pi})e^{\gamma C_{\pi}^{-1}}(C_{0}-C_{\pi})^{-1}(m_{0}-m_{\pi}).$ (3.4)
for the mean $b_{\gamma}$ and covariance $Q_{\gamma}$ of a single step of W-FR of size $\gamma$. Similarly, setting $M_{\gamma}=-\frac{1}{2}C_{\pi}(e^{2\gamma C_{\pi}^{-1}}-I)$ in Lemma B.2 yields
$P_{\gamma}$ $=C_{\pi}+\left(e^{\Gamma\gamma}(C_{0}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1}e^{2\gamma C_{\pi}^{-1}}\right)^{-1}$ (3.5)
$a_{\gamma}$ $=m_{\pi}+(P_{\gamma}-C_{\pi})e^{\gamma C_{\pi}^{-1}}(C_{0}-C_{\pi})^{-1}(m_{0}-m_{\pi}).$ (3.6)
for the mean $a_{\gamma}$ and covariance $P_{\gamma}$ of a single step of FR-W of size $\gamma$.
The structure of the mean for both splitting schemes is exactly the same as that obtained for the exact WFR flow with $C_{t}$ replaced by $Q_{\gamma},P_{\gamma}$ respectively. This demonstrates that in the Gaussian case the speed-up due to splitting is driven primarily through the improvement in covariance estimation. In particular, notice the difference in structure between $C_{t}$ vs $Q_{\gamma}$ and $P_{\gamma}$. The difference due to operator ordering is more subtle, as $Q_{\gamma}$ and $P_{\gamma}$ differ only by a single factor $e^{2\gamma C_{\pi}^{-1}}$. As will be shown later in this section, this factor does not necessarily lead to acceleration in all cases, and depends on the relationship between $\mu_{0}$ and $\pi$.
In order to compare the continuous evolution given by (3.1)-(3.2) with that given by the discrete time splitting schemes we consider $m_{n\gamma},C_{n\gamma}$, the mean and covariance respectively at time $t=n\gamma$ under the exact WFR flow. We will thus compare $m_{n\gamma},C_{n\gamma}$ with the mean and variance of the W-FR (and FR-W) split at step $n$ with step size $\gamma$.
Before presenting the main analytic results, consider Figure 3.1 which compares KL distance from $\pi$ after a single sequential split step and the same for the exact WFR in the univariate Gaussian setting. More specifically, it shows $\mathrm{KL}(\nu_{\gamma}||\pi)-\mathrm{KL}(\mu_{\gamma}||\pi)$ (blue line) and $\mathrm{KL}(\eta_{\gamma}||\pi)-\mathrm{KL}(\mu_{\gamma}||\pi)$ (red line) for a single sequential split step as a function of step size $\gamma$.
The KL is calculated exactly using the formula for Gaussians (B.24) and mean and covariance using the equations above. We consider the case where the target is more diffuse than the initial distribution (left plot, $C_{\pi}=100,m_{\pi}=20,C_{0}=1,m_{0}=0$) and the opposite case (right plot, $C_{\pi}=1,m_{\pi}=20,C_{0}=100,m_{0}=0$). Figure 3.1 shows that the improvement in KL due to splitting for larger $\gamma$ depends on the both the order of the operators and the on the relationship between the target and initial distribution. Specifically, in the univariate Gaussian case, when the target variance is larger than the initial variance, the ordering W-FR leads to a speed-up over the exact dynamics for all $\gamma>0$ see Figure 3.1 left panel. As expected for $\gamma$ large enough, the difference approaches zero as both the W and FR flows individually have been evaluated for a sufficiently long time so that the resulting densities are close to $\pi$. Conversely, when the target variance is smaller than the initial variance the ordering FR-W is superior, see Figure 3.1 right panel.
These numerical results are somewhat intuitive; if $\pi$ is more diffuse than $\mu_{0}$, applying the W operator first allows to increase variance and transform the initial distribution into a more diffuse one; if instead $\pi$ is more concentrated than $\mu_{0}$ applying the FR operator first allows to shrink the variance and returns a more concentrated distribution since the ratio $\pi/\mu_{0}$ will be small over most of the support. We will make this intuition more precise in the remainder of this section and extend to the multivariate Gaussian case, where we show that the difference in initial and target means plays a role also.
[…]
Figure 3.1: Difference in $\mathrm{KL}$ for a single time step $\gamma$ for W-FR split and FR-W on 1D Gaussians. Left: Target more diffuse than initial distribution ($m_{\pi}=20,C_{\pi}=100,m_{0}=0,C_{0}=1$). Right: Target more concentrated than initial distribution ($m_{\pi}=20,C_{\pi}=1,m_{0}=0,C_{0}=100$).
As a first step, we present an equality for the continuous time KL decay for the WFR PDE (2.1), which to the best of our knowledge is missing from the literature. Existing KL bounds for more general $\pi$ developed in the literature are not as tight as discussed in Section 5, see e.g. Figure 5.1(left).
The exact covariance and mean evolution in (3.4)-(3.3) for the W-FR split and (3.6)-(3.5) for the FR-W split can be used in a recursive fashion to characterise the decay of $\mathrm{KL}$ distance to $\pi$ of the split schemes compared to the exact WFR flow. We will make use of the following quantities throughout,
$E_{n}:=C_{n\gamma}-C_{\pi};\quad$ $E_{n}^{\beta}:=Q_{n\gamma}-C_{\pi};\quad E_{n}^{\alpha}:=P_{n\gamma}-C_{\pi};$ (3.7)
$\varepsilon_{n}:=m_{n\gamma}-m_{\pi};\quad$ $\varepsilon_{n}^{\beta}:=b_{n\gamma}-m_{\pi};\quad\varepsilon_{n}^{\alpha}:=a_{n\gamma}-m_{\pi}.$ (3.8)
Consider Figure 3.2 which shows the ratio of Split $\mathrm{KL}$ to exact WFR $\mathrm{KL}$ for a 10D Gaussian target (see Appendix B [us_neurips] for details) with $\gamma=0.7$. For this particular case, the W-FR scheme offers a speed-up over the exact dynamics, where $\mathrm{KL}$ from the split scheme is $\approx 60\%$ of $\mathrm{KL}$ in the exact case for large $t$. Conversely, a FR-W scheme leads to a reduction in convergence speed (see Figure 3.2 (right)). Importantly, while the ratio stabilises at approximately $1.8$, both $\mathrm{KL}(\eta_{n}||\pi)$ and $\mathrm{KL}(\mu_{n\gamma}||\pi)$ approach zero so that there is no asymptotic bias.
In the remainder of this section, we characterise KL decay and the properties giving rise to acceleration/deceleration. The following result characterises KL decay as a function of $n$ for the exact and sequential split schemes in the multivariate Gaussian case. The proof can be found in Appendix B.3.
Lemma 3.1.
KL decay, multivariate Gaussian. Consider the multivariate Gaussian setting, i.e. $\mu_{0}(x)=\mathcal{N}(x;m_{0},C_{0})$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$ with $C_{0},C_{\pi}$ strictly positive definite and $C_{0}\neq C_{\pi}$. Recall $\mu_{n\gamma}$ is the solution of (2.1) at time $t=n\gamma$ and $\nu_{n}$ is the solution of the W-FR split scheme at step $n$ with step size $\gamma$. Assume the exact WFR (2.1), W-FR split scheme (2.6) and FR-W split scheme (2.7) are initialised with the same $C_{0},m_{0}$ (i.e. $E_{0},\varepsilon_{0}$ fixed across methods), so that $\mu_{0}=\nu_{0}=\eta_{0}$. In particular, assume that $E_{0}$ is either strictly negative definite or strictly positive definite. Recall $\Gamma:=C_{\pi}^{-1}+\frac{1}{2}I$. Define also the functional
$\phi_{n}(A):=\frac{1}{2}\left[-\log\det(I+C_{\pi}^{-1}A)+Tr[C_{\pi}^{-1}A]+\varepsilon_{0}^{\top}E_{0}^{-1}e^{n\gamma C_{\pi}^{-1}}AC_{\pi}^{-1}Ae^{n\gamma C_{\pi}^{-1}}E_{0}^{-1}\varepsilon_{0}\right],$
where $A$ denotes any $d\times d$ symmetric matrix and the matrix operator
$J_{n}(B)=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+B\,C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\right)^{-1}e^{-n\gamma\Gamma},$
for any $B\in\mathbb{R}^{d\times d}$. Then the following holds for all $n=1,2,3,\dots$,
  1. 1.
    The solution of the exact WFR PDE (2.1) at time $t=n\gamma$ satisfies
    $\mathrm{KL}(\mu_{n\gamma}||\pi)$ $=\phi_{n}(E_{n}),$ (3.9)
    $E_{n}$ $=J_{n}(\Omega),$ (3.10)
    $\Omega$ $:=\frac{1}{2}\Gamma^{-1}.$ (3.11)
  2. 2.
    The recursive solution of the W-FR split PDE in Proposition 2.2 after $n$ time steps of size $\gamma$ satisfies
    $\mathrm{KL}(\nu_{n}||\pi)$ $=\phi_{n}(E_{n}^{\beta}),$ (3.12)
    $E_{n}^{\beta}$ $=J_{n}(\Omega^{\beta}),$ (3.13)
    $\Omega^{\beta}$ $:=(1-e^{\gamma})(I-e^{2\gamma\Gamma})^{-1}.$ (3.14)
  3. 3.
    The recursive solution of the FR-W split PDE in Proposition 2.3 after $n$ time steps of size $\gamma$ satisfies
    $\mathrm{KL}(\eta_{n\gamma}||\pi)$ $=\phi_{n}(E_{n}^{\alpha}),$ (3.15)
    $E_{n}^{\alpha}$ $=J_{n}(\Omega^{\alpha}),$ (3.16)
    $\Omega^{\alpha}$ $:=(1-e^{\gamma})(I-e^{2\gamma\Gamma})^{-1}e^{2\gamma C_{\pi}^{-1}}=(1-e^{\gamma})(e^{-2\gamma C_{\pi}^{-1}}-e^{\gamma})^{-1}.$ (3.17)
Lemma 3.1 shows that the change in KL decay due to splitting is, as discussed previously, entirely due to differences in covariance estimation. Despite the considerable difference in structure of $C_{\gamma}$ vs $Q_{\gamma}$ or $P_{\gamma}$ (c.f. (3.1) with $t=\gamma$ and (3.3), (3.5)), the $n$-step iterations lead to expressions that more clearly highlight similarities between the exact and sequential splitting. More specifically, notice that the difference quantities $E_{n}$ and $E_{n}^{\beta},E_{n}^{\alpha}$ are identical up to the $\Omega$ and $\Omega^{\beta},\Omega^{\alpha}$ factors.
This brings us to our main result of this section, the proof of which can be found in Appendix B.4. It quantifies acceleration or deceleration over the exact WFR dynamics due to splitting in terms of KL decay.
Proposition 3.2.
Convergence speed, multivariate Gaussian. Assume the same conditions as in Lemma 3.1. Then the following holds:
  1. (i)
    Suppose $E_{0}\succ 0$. Then $E_{n}\succ 0,E_{n}^{\beta}\succ 0$ and $E_{n}^{\alpha}\succ 0$ for all $n=1,2,3\dots$.
  2. (ii)
    Suppose $E_{0}^{-1}C_{\pi}\prec-\Omega$, $E_{0}^{-1}C_{\pi}\prec-\Omega^{\beta}$ and $E_{0}^{-1}C_{\pi}\prec-\Omega^{\alpha}$. Then $E_{n}\prec 0,E_{n}^{\beta}\prec 0$ and $E_{n}^{\alpha}\prec 0$ for all $n=1,2,3,\dots$.
  3. (iii)
    Assume the conditions of (i) or (ii). Then the asymptotic (large $n$) ratio of the $\mathrm{KL}$ distance to $\pi$ of the W-FR split scheme solution with respect the solution of the exact WFR is given by
    $\lim_{n\rightarrow\infty}\frac{\mathrm{KL}(\nu_{n}||\pi)}{\mathrm{KL}(\mu_{n\gamma}||\pi)}=\frac{p_{1}^{\top}(E_{0}^{-1}C_{\pi}+\Omega^{\beta})^{-1}D_{0}(E_{0}^{-1}C_{\pi}+\Omega^{\beta})^{-1}p_{1}}{p_{1}^{\top}(E_{0}^{-1}C_{\pi}+\Omega)^{-1}D_{0}(E_{0}^{-1}C_{\pi}+\Omega)^{-1}p_{1}},$ (3.18)
    where $\Omega$ and $\Omega^{\beta}$ are as defined in (3.11) and (3.14) respectively, $p_{1}$ is the eigenvector corresponding to the smallest eigenvalue of $\Gamma$ and $D_{0}:=E_{0}^{-1}\varepsilon_{0}\varepsilon_{0}^{\top}E_{0}^{-1}C_{\pi}$. Likewise, for the FR-W split scheme, the same limiting ratio for $\frac{\mathrm{KL}(\eta_{n\gamma}||\pi)}{\mathrm{KL}(\mu_{n\gamma}||\pi)}$ holds with $\Omega^{\beta}$ replaced by $\Omega^{\alpha}$ as defined in (3.17).
Notice that in dimension 1, the limiting ratio in (3.18) is independent of $\varepsilon_{0}$, so that the speed-up depends only on the initial covariance relative to the target covariance. In particular, in the 1-dimensional case if $C_{0}>C_{\pi}$ then FR-W will converge faster than the exact WFR dynamics, conversely, if $C_{\pi}>C_{0}$ then W-FR will be faster. In the multivariate case however, the squared mean difference $\varepsilon\varepsilon^{\top}$ plays an important role also.
[…]
Figure 3.2: Left: Ratio of KL from $n$-step W-FR scheme to KL from exact WFR as a function of $t=n\times\gamma$, $\pi$ is a 10D Gaussian (see Appendix B [us_neurips] for details) and $\gamma=0.7$. The horizontal black line corresponds to (3.18). For reference, $\mathrm{KL}(\mu_{4.2}||\pi)=9.7$. Right: Same as left but for FR-W scheme.
4 Preservation of log-concavity of exact WFR
The previous section shows that, in the Gaussian case, splitting schemes can speed up convergence provided $\mu_{0}$ is chosen wisely.
In order to generalise this result to say, log-concave target distributions, we first provide a result demonstrating that log-concavity is preserved under the evolution of the exact WFR flow (2.1). Recall that for any log-concave $\mu_{0}$, preservation of log-concavity uniformly in time under the W flow is guaranteed only for $\pi$ Gaussian (i.e. for Ornstein-Uhlenbeck Semigroups) due to [Kolesnikov2001]. This is in stark contrast to the FR flow, where log-concavity is preserved uniformly in time for any strongly log-concave $\pi$ and $\mu_{0}$ (see Lemma C.1).
With some stronger conditions on $\mu_{0}$ and $\pi$, as detailed in Assumptions 2 and 3, we establish preservation of log-concavity under the W flow for a limited time horizon, using similar arguments as in [Pathiraja2021, Section 7.2] (in the context of stochastic filtering) and Lemma 16 in [liang_characterizing_2025]. We first state the necessary assumptions:
Assumption 2.
In addition to Assumption 1, the following conditions hold
  1. (a)
    $\pi(x)\propto e^{-V_{\pi}(x)}$ is $\alpha_{\pi}$-strongly log-concave and $L_{\pi}$-smooth, i.e. there exists an $L_{\pi}\geq\alpha_{\pi}>0$ such that $L_{\pi}I\succeq\nabla^{2}V_{\pi}(x)\succeq\alpha_{\pi}I$ for all $x\in\mathbb{R}^{d}$;
  2. (b)
    $\mu_{0}(x)\propto e^{-V_{0}(x)}$ is $\alpha_{0}$-strongly log-concave and $L_{0}$-smooth, i.e. there exists an $L_{0}\geq\alpha_{0}>0$ such that $L_{0}I\succeq\nabla^{2}V_{0}(x)\succeq\alpha_{0}I$ for all $x\in\mathbb{R}^{d}$.
Assumption 3.
Given Assumption 2, the following conditions hold
  1. (a)
    $V_{0}-\frac{(1+\delta)}{2}V_{\pi}$ is strongly convex with parameter $\alpha_{d}>0$ for some specified $0<\delta<1$.
  2. (b)
    Denote by $R:=-\Delta V_{\pi}+|\nabla V_{\pi}|^{2}$. Also define $\mathcal{H}:=R+V_{\pi}$. Assume $\mathcal{H}$ is strongly convex with parameter $\alpha_{h}>0$. This implies that
    $\nabla^{2}R\succeq(\alpha_{h}-L_{\pi})I,$
    where $\alpha_{h}-L_{\pi}<0$.
The next lemma gives the precise statement of the time horizon over which the W flow is guaranteed to preserve strong log-concavity. The proof can be found in Appendix C.1.
Lemma 4.1.
W flows preserve log-concavity over a fixed time horizon. Suppose $\mu_{t}$ is the solution of the W flow (2.2) at time $t$, initialised at $\mu_{0}(x)$ satisfying Assumption 3. Let $b:=\sqrt{|\alpha_{h}-L_{\pi}|}$. Then there exists a $t^{\ast}>0$ such that for all $t<t^{\ast}$, $\mu_{t}(x)\propto e^{-\mathcal{E}_{t}(x)}$, with $\mathcal{E}_{t}(x)$ a strongly convex function,
$\nabla^{2}\mathcal{E}_{t}(x)$ $\succeq\left(\frac{\alpha_{\pi}}{2}+c_{t}\right)I,\quad\forall\enskip x\in\mathbb{R}^{d},$ (4.1)
$c_{t}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-bt\right)>0,$ (4.2)
$c_{0}$ $=\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}.$ (4.3)
Firstly, notice that in Lemma 4.1, we consider the case where $\alpha_{h}-L_{\pi}<0$, since $\alpha_{h}>L_{\pi}$ implies $R$ is strongly convex. Strong convexity of $R$ would imply uniform in time convexity of $\mu_{t}$ by similar techniques as in the proof of Lemma 4.1, which is not guaranteed unless $\pi$ is Gaussian due to [Kolesnikov2001]. The length of the time horizon over which strong convexity is preserved also depends on the log-concavity of the difference between initial and target potentials.
In the next theorem, we show that the preservation of log-concavity under the FR flow can be exploited to ensure the same for the WFR flow, uniformly in time. The proof of the theorem can be found in Appendix C.2.
Theorem 4.1.
WFR preserves log-concavity uniformly in time. Assume the conditions of Lemma 4.1 and that $b:=\sqrt{\frac{|\alpha_{h}-L_{\pi}|}{2}}$ satisfies
$b^{2}<\tfrac{\alpha_{\pi}}{2}.$ (4.4)
Then $\mu_{t}$, the solution of (2.1) at time $t$ is $\alpha_{t}$-strongly log-concave for all $t>0$, where
$\alpha_{t}=\frac{\alpha_{\pi}}{2}+\frac{c_{\infty,1}-c_{\infty,2}l_{0}e^{-t(c_{\infty,1}-c_{\infty,2})}}{1-l_{0}e^{-t(c_{\infty,1}-c_{\infty,2})}},$ (4.5)
and
$l_{0}$ $=\frac{(c_{0}-c_{\infty,1})}{(c_{0}-c_{\infty,2})};\qquad c_{\infty,1}=\frac{-1+\sqrt{1+4(\frac{\alpha_{\pi}}{2}-b^{2})}}{2};\qquad c_{\infty,2}=\frac{-1-\sqrt{1+4(\frac{\alpha_{\pi}}{2}-b^{2})}}{2};$
and $c_{0}$ is given by (4.3).
The tightness of the constants (4.5) is evaluated empirically for the 1D Gaussian case in Figure 4.1 where the true constants are known. As can be seen there, for a sufficiently large difference in convexity of the initial and target potentials, the constants (4.5) are reasonably tight (c.f. left vs right). We will rely on the results of Theorem 4.1 in the next section to characterise $\mathrm{KL}$ convergence rates.
Figure 4.1: Comparison of the log-concavity constant predicted by Theorem 4.1 and the true log-concavity constant obtained from (3.1) for a 1D Gaussian with $\mu_{0}(x)=\mathcal{N}(x;,0,1)$ and $\pi(x)=\mathcal{N}(x;,0,100)$ (left), $\pi(x)=\mathcal{N}(x;,0,5)$ (middle), $\pi(x)=\mathcal{N}(x;,0,2.1)$ (right). For $C_{\pi}\leq 2$ the assumptions of Theorem 4.1 are not satisfied.
5 Convergence speed of Exact WFR: log-concave case
The rate of decay of $\mathrm{KL}$ for the WFR flow is obtained considering the time derivative of $\mathrm{KL}(\mu_{t}||\pi)$ along the WFR flow
$\frac{d}{dt}\mathrm{KL}(\mu_{t}||\pi)$ $=\int\log\frac{\mu_{t}}{\pi}\partial_{t}\mu_{t}=-\int\mu_{t}|\nabla\log\frac{\mu_{t}}{\pi}|^{2}-\text{Var}_{\mu_{t}}\left[\log\frac{\mu_{t}}{\pi}\right].$ (5.1)
As observed in [gallouet2017jko, page 13], the first term corresponds to the negative gradient of $\mathrm{KL}$ w.r.t. the Wasserstein-2 metric, while the remaining terms give the negative gradient of $\mathrm{KL}$ w.r.t. the Fisher–Rao geometry, implying that the dissipation for the WFR metric is the sum of the W and the FR dissipation. As a consequence, the convergence rate of WFR gradient flow (2.1) can be upper bounded by
$\mathrm{KL}(\mu_{t}||\pi)\leq\min\left\{\mathrm{KL}(\mu_{t}^{\textrm{FR}}|\pi),\mathrm{KL}(\mu_{t}^{\textrm{W}}|\pi)\right\},$ (5.2)
where $\mu_{t}^{\textrm{FR}},\mu_{t}^{\textrm{W}}$ denote the solutions of the FR and W flow PDEs respectively. A bound on the decay of KL along the W flow can be obtained if $\pi$ satifies an LSI (2.8) (e.g., [chewi2024analysis]). Decay of KL along the FR flow at rate $\mathcal{O}(e^{-t})$ is guaranteed under boundedness conditions on the score $\log\pi/\mu_{0}$ ([chen2023sampling, Theorem 4.1] and [lu2023birth, Proposition 2.4]).
A sharper rate of converge for WFR can be obtained assuming $\log\nicefrac{{\pi(x)}}{{\mu_{0}(x)}}\geq M$ and a warm-start condition [Lu2019, Appendix B]
$\mathrm{KL}(\mu_{t}||\pi)$ $\leq e^{-\left(2\lambda_{\pi}^{-1}+(2-3\delta)\right)(t-t_{0})}\mathrm{KL}(\mu_{0}||\pi),$ (5.3)
for all $t\geq t_{0}:=\log(M/\delta^{3})$ and $\delta>0$.
We point out that neither (5.2) nor (5.3) provide a sharp rate of decay for $\mathrm{KL}(\mu_{t}||\pi)$. In fact, (5.2) shows that the rate of WFR is never worse than that of W or FR, looking at (5.1) however it is evident that the rate of decay of the WFR flow should be considerably higher than that of the two singular flows. Similarly, (5.3) shows that the decay in the case of the WFR flow is faster than that of the W flow but the rate depends on an arbitrary constant $\delta$ which makes (5.3) harder to interpret.
In Proposition 5.1, we obtain a result that shows the decay rate is indeed the sum of the W and FR rates, but for the symmetrised KL, not KL. The symmetrised $\mathrm{KL}$ divergence, or Jeffrey’s divergence $J$, is defined as
$J(\mu,\pi):=\mathrm{KL}(\mu||\pi)+\mathrm{KL}(\pi||\mu),$
for any $\mu\ll\pi$ and also $\pi\ll\mu$ . The rationale behind considering the symmetrised KL is due to the fact that the KL is not geodesically convex under the FR flow [carrillo_fisher-rao_2024, Theorem 1.1] nor satisfies a gradient dominance condition [carrillo_fisher-rao_2024, Theorem 4.1], which makes its convergence analysis more difficult. On the other hand, Jeffrey’s divergence satisfies a gradient dominance condition [carrillo_fisher-rao_2024, Section 4] which allows to achieve improved rates. In addition, $J(\mu,\pi)$ upper bounds $\mathrm{KL}(\mu||\pi)$ and thus the obtained rates for $J(\mu,\pi)$ gives information on the decay of KL too. The Kl decay rate obtained in Proposition 5.1 is a sum of the decay rates of W and FR flow, conjectured by [domingo-enrich2023an] but not yet proven in the literature.
Theorem 4.1 immediately provides sufficient conditions to establish an exponential decay result of the symmetrised $\mathrm{KL}$ divergence, or Jeffrey’s divergence $J$. The proof makes use of recent functional inequality results on pure FR flows [carrillo_fisher-rao_2024, Section 4] and provides a sharper rate of convergence than the results available in the literature (Figure 5.1).
Proposition 5.1.
Let Assumption 2 hold. Then the following decay result holds for $\mu_{t}$ the solution of the WFR PDE (2.1) for all $t>0$
$J(\mu_{t},\pi)\leq J(\mu_{0},\pi)e^{-t\kappa},$
where $\kappa=2\min(\alpha_{\pi},\alpha_{t})+1$ and $\alpha_{t}$ is given in Theorem 4.1.
Proof.
Assuming valid interchange of differentiation and integration,
$\frac{d}{dt}J(\mu_{t},\pi)$ $=\int\left(\log\frac{\mu_{t}}{\pi}-\frac{\pi}{\mu_{t}}\right)(x)\partial_{t}\mu_{t}(x)dx$
$=\int\left(\log\frac{\mu_{t}}{\pi}-\frac{\pi}{\mu_{t}}\right)(x)\left(f_{\textrm{W}}(\mu_{t}(x))+f_{\textrm{FR}}(\mu_{t}(x))\right)dx.$ (5.4)
Begin with the first term, by a standard integration by parts argument,
$\int\left(\log\frac{\mu_{t}}{\pi}-\frac{\pi}{\mu_{t}}\right)f_{\textrm{W}}(\mu_{t})dx$ $=-\int\left(\log\frac{\mu_{t}(x)}{\pi(x)}-\frac{\pi(x)}{\mu_{t}(x)}\right)\nabla\cdot\left(\mu_{t}(x)\nabla\log\frac{\mu_{t}}{\pi}(x)\right)(x)dx$ (5.5)
$=-\mathbb{E}_{\mu_{t}}\left[\left|\nabla\log\frac{\mu_{t}}{\pi}\right|^{2}\right]-\mathbb{E}_{\pi}\left[\nabla\log\frac{\mu_{t}}{\pi}\cdot\nabla\log\frac{\mu_{t}}{\pi}\right]$
$=-(\mathcal{I}(\mu_{t}||\pi)+\mathcal{I}(\pi||\mu_{t}))$
$\leq-2\min(\lambda_{\pi}^{-1},\lambda_{\mu_{t}}^{-1})(\mathrm{KL}(\mu_{t}||\pi)+\mathrm{KL}(\pi||\mu_{t}))$
$=-2\min(\alpha_{\pi},\alpha_{t})J(\mu_{t},\pi),$
where $\mathcal{I}(\mu_{t}||\pi)$ is the relative Fisher information between $\mu_{t}$ and $\pi$. The last inequality holds due to Assumption 2 and Theorem 4.1. More specifically, since $\pi$ is $\alpha_{\pi}$-strongly log-concave, it immediately satisfies a Log-Sobolev inequality (2.8) with constant $\lambda_{\pi}=\alpha_{\pi}^{-1}$ (likewise for $\mu_{t}$, which is $\alpha_{t}$-strongly log-concave due to Theorem 4.1). Then for the second term in (5.4), by a direct application of (5.16) in [carrillo_fisher-rao_2024, Theorem 5.6] with $f(y)=y\log(y)$ and $\bar{f}=yf(y^{-1})$, it holds that
$-\int\frac{\pi}{\mu_{t}}f_{\textrm{FR}}(\mu_{t})dx\leq-J(\mu_{t},\pi)l,$
and also,
$\int\log\frac{\mu_{t}}{\pi}f_{\textrm{FR}}(\mu_{t})dx=-\text{Var}_{\mu_{t}}\left(\log\frac{\mu_{t}}{\pi}\right)<0,$
so that together, we have
$\int\left(\log\frac{\mu_{t}}{\pi}-\frac{\pi}{\mu_{t}}\right)f_{\textrm{FR}}(\mu_{t})dx\leq-J(\mu_{t},\pi).$ (5.6)
An application of Grönwall’s lemma then yields the result. ∎
Refer to captionRefer to caption
Figure 5.1: Comparison of exact KL decay (left) and symmetrised KL decay (right) for a 1D Gaussian with $m_{\pi}=20,C_{\pi}=100,m_{0}=0,C_{0}=1$. Left plot: We compare the exact decay of KL from Proposition 3.2 with the rates in (5.2) and (5.3) with $\delta=0.1,t_{0}=6.9$. Right plot: We compare the exact decay of the symmetrised KL obtained using the mean and covariance evolution in (3.1)–(3.2) with the rate in Proposition 5.1.
6 Splitting & convergence speed
Existing $\mathrm{KL}$ decay results for the W flow and FR flow developed in the literature cannot be straightforwardly utilised to characterise differences in convergence speed between the W-FR and FR-W scheme. In particular, combining existing exponential decay bounds for the KL divergence for the W part and the FR part would not allow us to isolate changes to convergence speed depending on the order as the exponents simply add (and commute). Hence, the first variation results developed in Proposition 2.2 and 2.3 will be critical in characterising how convergence speed is affected due to splitting, and the ordering of operators.
We utilise the variational formulae (2.9) and (2.14) to tighten the upper bound on the KL. We focus on one single split step of size $\gamma>0$ and consider the classical KL divergence $\mathrm{KL}(\mu||\pi)$, as a first step.
6.1 Convergence speed: W-FR
We first focus on the PDE (2.9) due to its easier form. It is possible to obtain a KL decay ODE by a straightforward calculation using (2.9), integration by parts, the sufficient tail decay of $\nu_{\gamma}$ and $\pi$ guaranteed by Assumption 1 and assuming sufficient regularity to interchange differentiation and integration.
Consider one step of the W-FR splitting (2.6) of size $\gamma>0$. Recall the split W-FR PDE in Proposition 2.2
$\partial_{\gamma}\nu_{\gamma}=f_{\textrm{W}}(\nu_{\gamma})+f_{\textrm{FR}}(\nu_{\gamma})+(e^{\gamma}-1)f_{\textrm{P}}(\nu_{\gamma}),$
where $f_{\textrm{P}}$ is in (2.12). Then, we have
$\frac{d}{d\gamma}\mathrm{KL}(\nu_{\gamma}||\pi)$ $=\int\log\frac{\nu_{\gamma}(x)}{\pi(x)}\left(f_{\textrm{W}}(\nu_{\gamma}(x))+f_{\textrm{FR}}(\nu_{\gamma}(x))+(e^{\gamma}-1)f_{\text{P}}(\nu_{\gamma}(x))\right)dx$ (6.1)
$=-\mathcal{I}(\nu_{\gamma}||\pi)-\text{Var}_{\nu_{\gamma}}\left[\log\frac{\nu_{\gamma}}{\pi}\right]+(e^{\gamma}-1)\text{Cov}_{\nu_{\gamma}}\left(\log\frac{\nu_{\gamma}}{\pi},\left|\nabla\log\frac{\nu_{\gamma}}{\pi}\right|^{2}\right),$
where $\mathcal{I}(\nu_{\gamma}||\pi)$ is the relative Fisher information.
It is clear that the behaviour of KL decay for the W-FR split will depend on the perturbation term $\text{Cov}_{\nu_{\gamma}}\left(g(\nu_{\gamma}),\left|\nabla g(\nu_{\gamma})\right|^{2}\right)$. This quantity measures the covariation between the log density ratio (when it is pointwise positive, it means that the current density is larger than the target density) and the squared score function (or magnitude of the score function), which is large in regions where the ratio $\frac{\nu_{\gamma}(x)}{\pi(x)}$ varies rapidly. In fact, if the log-ratio and the score are strongly negatively correlated, it is this negative correlation that contributes to $\text{Cov}_{\nu_{\gamma}}\left(g(\nu_{\gamma}),\left|\nabla g(\nu_{\gamma})\right|^{2}\right)$ being negative and driving the speed-up. In the below Lemma, we list conditions under which this covariance is negative. We emphasise that these conditions are sufficient, and establishing similar results for more relaxed conditions is left for future work. Given that a similar speed-up can be achieved for a 1D Gaussian mixture target (see Figure 1 (right) in [us_neurips]), we expect this to be possible.
However, we emphasise that negativity of the perturbation term is not sufficient to guarantee a speed-up, uniformly in time. A more precise characterisation, e.g. via comparison theorems, would require a lower-bound on the decay rate of the $\mathrm{KL}$. This aspect is left for future work.
Lemma 6.1.
The following results hold.
  1. (i)
    [us_neurips, Proposition 2] Consider $\nu(x)=\mathcal{N}(x;b,Q)$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$. If $Q\prec C_{\pi}$ then $\text{Cov}_{\nu}\left(g(\nu),\left|\nabla g(\nu)\right|^{2}\right)<0$.
  2. (ii)
    Denote $h(x):=-g(\nu(x))$. Assume $h$ is symmetric and $\nabla^{2}h\succeq\alpha I$. Then $\text{Cov}_{\nu}\left(g(\nu),\left|\nabla g(\nu)\right|^{2}\right)<0$.
Proof.
We assume w.l.o.g. that $h$ is symmetric about 0 and that $h\geq 0$ since it is strongly convex and thus lower bounded. Since $h$ and $|\nabla h|^{2}$ are both decreasing in - and both increasing in + we can apply Chebyshev’s integral inequality (e.g. [egozcue2009some]) to each interval and obtain
$\int\nu(x)h(x)|\nabla h(x)|^{2}dx$ $=\frac{1}{\nu(_{-})}\int_{{}_{-}}\nu(x)h(x)|\nabla h(x)|^{2}dx+\frac{1}{\nu(_{+})}\int_{{}_{+}}\nu(x)h(x)|\nabla h(x)|^{2}dx$
$\geq\frac{1}{\nu(_{-})^{2}}\int_{{}_{-}}\nu(x)h(x)dx\int_{{}_{-}}\nu(x)|\nabla h(x)|^{2}dx$
$+\frac{1}{\nu(_{+})^{2}}\int_{{}_{+}}\nu(x)h(x)dx\int_{{}_{+}}\nu(x)|\nabla h(x)|^{2}dx.$
Due to the symmetry of $h$ we have $\int_{{}_{+}}\nu(x)h(x)dx=\int_{{}_{-}}\nu(x)h(x)dx$ and similarly for $|\nabla h|^{2}$. Using the fact that $\nu(_{+}),\nu(_{-})\leq 1$ since $\nu\in\mathcal{P}_{2}^{ac}(^{d})$, we finally have
$\int\nu(x)h(x)|\nabla h(x)|^{2}dx$ $\geq\int\nu(x)h(x)dx\int\nu(x)|\nabla h(x)|^{2}dx.$
Recalling that g=hg=-h, the result follows. ∎
Inequalities of the form $\text{Cov}_{\nu}(\alpha(X),\beta(X))<0$ for general measures $\mu$ and functions $\alpha,\beta$ are an active area of research [egozcue2011covariance, bonnefont2024covariance, egozcue2009some, saumard2014log] and we believe that bounds obtained from (6.1) can inform the choice of splitting to choose for a given pair of target $\pi$ and initial distribution $\mu_{0}$ provided these inequalities are obtained.
6.2 Convergence speed: FR-W
Similarly to the W-FR we can use the split PDE (2.14) to quantify the differen in KL decay due to splitting over one step of size $\gamma>0$.
Consider one step of the FR-W splitting (2.7) of size $\gamma>0$. We have
$\frac{d}{d\gamma}\mathrm{KL}(\eta_{\gamma}||\pi)$ $=\int\log\frac{\eta_{\gamma}(x)}{\pi(x)}\left(f_{\textrm{W}}(\eta_{\gamma}(x))+f_{\textrm{FR}}(\eta_{\gamma}(x))+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,-f_{\textrm{p}}(\eta_{\tau})\right)(x)d\tau\right)dx$
$=-\mathcal{I}(\eta_{\gamma}||\pi)-\text{Var}_{\eta_{\gamma}}\left[\log\frac{\eta_{\gamma}}{\pi}\right]-\int_{0}^{\gamma}\text{Cov}_{\eta_{\tau}}\left(S_{\textrm{W}}\left(\gamma-\tau,g(\eta_{\gamma})\right),|\nabla g(\eta_{\tau})|^{2}\right)d\tau.$
In the below Lemma, we list conditions under which
$\int_{0}^{\gamma}\text{Cov}_{\eta_{\tau}}\left(S_{\textrm{W}}\left(\gamma-\tau,g(\eta_{\gamma})\right),|\nabla g(\eta_{\tau})|^{2}\right)d\tau>0.$ (6.2)
Lemma 6.2.
Let $g(\eta_{\tau}(x))$ be symmetric with $\nabla^{2}g(\eta_{\tau})\succeq\alpha_{\tau}I$ for all $\tau\in[0,\gamma]$. Then (6.2) holds.
Proof.
We assume w.l.o.g. that $g(\eta_{\tau})$ is symmetric about 0 and that $g(\eta_{\tau})\geq 0$ since $g(\eta_{\tau})$ is strongly convex and thus lower bounded. Recall that
$S_{\textrm{W}}(t,v)(x)$ $=\int(4\uppi t)^{-d/2}\exp\left(-\frac{|x-y-\int_{0}^{t}\nabla\log\pi(X_{s})ds|^{2}}{4t}\right)v(y)dy,$
where $(X_{t})_{t\geq 0}$ solves $dX_{t}=\nabla\log\pi(X_{t})dt+\sqrt{2}dW_{t}$. It follows that $S_{\textrm{W}}\left(\gamma-\tau,g(\eta_{\gamma})\right)\geq 0$ is also symmetric about 0. Proceeding as in the proof of Lemma 6.1 gives $\text{Cov}_{\eta_{\tau}}\left(S_{\textrm{W}}\left(\gamma-\tau,g(\eta_{\gamma})\right),|\nabla g(\eta_{\tau})|^{2}\right)>0$ from which the result follows.
7 Conclusion & Future Work
Operator splitting methods lie at the heart of many gradient flow based algorithms, in particular, for discretisations of the Wasserstein–Fisher–Rao gradient flow. We examine errors introduced by splitting schemes whereby the Wasserstein and Fisher–Rao operators are solved separately (and exactly). We develop variational formulae for the split schemes and identify a perturbation which modifies the dynamics. It is demonstrated that the order in which these operators are evaluated can have a considerable impact on convergence speed, without affecting $\pi$-invariance. This impact is unsurprisingly more pronounced when the chosen step size is large, but not so large that it is nearly equivalent to evaluating either the Wasserstein or Fisher-Rao flow alone. In practical terms, this means that a sequence of alternating operators with a fixed step size $\gamma$ can be re-arranged (e.g. such that $m$ consecutive steps of size $\gamma$ of the same operator are evaluated) in order to reduce time to convergence, with no extra computational burden.
We focus primarily on characterising this behaviour in the Gaussian case and then more broadly in the strongly log-concave case, showing that although the exact Wasserstein–Fisher–Rao flow preserves log-concavity uniformly in time, under some conditions. Decay rates of the exact WFR flow are then characterised in terms of the symmetrised $\mathrm{KL}$, using the derived log-concavity constants and new functional inequalities for Fisher–Rao flows in [carrillo_fisher-rao_2024].
There are several avenues for future work, including (but not limited to) 1) quantifying the impact of further space and time discretisations of the W and FR flows, which are currently assumed to be evaluated exactly; 2) using the derived results to develop adaptive step size splitting schemes and to determine which ordering is preferrable and finally 3) in more precisely quantifying the acceleration/deceleration of symmetrised $\mathrm{KL}$ due to splitting.
Acknowledgements
The authors would like to thank Josh Bon, Sam Power, Florian Maire, Daniel Paulin, Upanshu Sharma for helpful discussions and in particular Sinho Chewi for pointing to reference [Kolesnikov2001] and Andre Wibisono for directing us to Lemma 16 in [liang_characterizing_2025].
Appendix A Auxiliary lemmas
Lemma A.1.
Let Assumption 1 hold. Recall the $k$th commutator $\mathcal{G}^{k}\eta_{\gamma}:=[f_{\textrm{W}},\mathcal{G}^{k-1}\eta_{\gamma}]$ for $k=1,2,\dots$ with $\mathcal{G}^{0}:=f_{\textrm{FR}}$ and $\mathcal{G}^{1}$ as defined in (2.17). The following identity holds
$\int_{0}^{\gamma}e^{(\gamma-\tau)f_{\textrm{W}}}\mathcal{G}^{1}\eta_{\tau}d\tau=\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}^{k}(\eta_{\gamma}).$
Proof.
Define $f(\tau)=e^{(\gamma-\tau)f_{W}}\mathcal{G}^{1}\eta_{\tau}$ and consider a Taylor expansion around point $\gamma$ of $f$, $f(\tau)=\sum_{k=0}^{\infty}\frac{f^{(k)}(\gamma)}{k!}(\tau-\gamma)^{k}$. Let us compute the derivatives of $f$. Applying the chain rule we have
$\partial_{s}e^{(\gamma-s)f_{\textrm{W}}}\mathcal{G}^{1}\eta_{s}$ $=-f_{\textrm{W}}e^{(\gamma-s)f_{\textrm{W}}}\mathcal{G}^{1}\eta_{s}+e^{(\gamma-s)f_{\textrm{W}}}\partial_{s}\mathcal{G}^{1}\eta_{s}$
Recall that
$\mathcal{G}^{1}v=[f_{\textrm{W}},f_{\textrm{FR}}](v)=f_{\textrm{W}}(f_{\textrm{FR}}(v))-f_{\textrm{FR}}^{\prime}(v)[f_{\textrm{W}}(v)].$
By chain rule and linearity of $f_{\textrm{W}}$,
$\partial_{s}f_{\textrm{W}}(f_{\textrm{FR}}(\eta_{s}))=f_{\textrm{W}}^{\prime}(f_{\textrm{FR}}(\eta_{s}))\left[\partial_{s}f_{\textrm{FR}}(\eta_{s})\right]=f_{\textrm{W}}(f^{\prime}_{\textrm{FR}}(\eta_{s})[\partial_{s}\eta_{s}])=f_{\textrm{W}}(f^{\prime}_{\textrm{FR}}(\eta_{s})[f_{\textrm{W}}(\eta_{s})])$ (A.1)
since $\partial_{s}\eta_{s}=\partial_{s}S_{\textrm{W}}(s,u_{\gamma})=f_{\textrm{W}}(\eta_{s})$. To compute $\partial_{s}f_{\textrm{FR}}^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]$ we recall that
$\partial_{s}f_{\textrm{FR}}^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]$ $=f_{\textrm{FR}}^{\prime\prime}(\eta_{s})(\partial_{s}\eta_{s},f_{\textrm{W}}(\eta_{s}))+f_{\textrm{FR}}^{\prime}(\eta_{s})[\partial_{s}f_{\textrm{W}}(\eta_{s})]$ (A.2)
$=f_{\textrm{FR}}^{\prime\prime}(\eta_{s})(f_{\textrm{W}}(\eta_{s}),f_{\textrm{W}}(\eta_{s}))+f_{\textrm{FR}}^{\prime}(\eta_{s})[\partial_{s}f_{\textrm{W}}(\eta_{s})]$
$=f_{\textrm{FR}}^{\prime\prime}(\eta_{s})(f_{\textrm{W}}(\eta_{s}),f_{\textrm{W}}(\eta_{s}))+f_{\textrm{FR}}^{\prime}(\eta_{s})[f_{\textrm{W}}^{2}(\eta_{s})]$
where in the last line we have used $\partial_{s}\eta_{s}=f_{\textrm{W}}(\eta_{s})$ and $\partial_{s}^{2}\eta_{s}=\partial_{s}f_{\textrm{W}}(\eta_{s})=f_{\textrm{W}}^{\prime}(\eta_{s})[\partial_{s}\eta_{s}]=f_{\textrm{W}}(f_{\textrm{W}}(\eta_{s}))=f_{\textrm{W}}^{2}(\eta_{s})$ by chain rule, and where $f_{\textrm{FR}}^{\prime\prime}(\eta_{s})[\partial_{s}\eta_{s},f_{\textrm{W}}(\eta_{s})]$ is a bilinear map describing how the directional derivative $f_{\textrm{FR}}^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]$ changes in the direction $\partial_{s}\eta_{s}$.
Combining (A.1)–(A.2) it the immediately follows that
$\partial_{s}\mathcal{G}^{1}\eta_{s}$ $=f_{\textrm{W}}(f^{\prime}_{\textrm{FR}}(\eta_{s})[f_{\textrm{W}}(\eta_{s})])-\left(f_{\textrm{FR}}^{\prime\prime}(\eta_{s})(f_{\textrm{W}}(\eta_{s}),f_{\textrm{W}}(\eta_{s}))+f_{\textrm{FR}}^{\prime}(\eta_{s})[f_{\textrm{W}}^{2}(\eta_{s})]\right)$
$=f_{\textrm{W}}(f_{\textrm{FR}}^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})])-(f_{\textrm{FR}}^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})])^{\prime}[f_{\textrm{W}}(\eta_{s})]$
$=[f_{\textrm{W}},f_{\textrm{FR}}]^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]$
$=(\mathcal{G}^{1})^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]$
using the chain rule for operators
$(g^{\prime}(v)[f(v)])^{\prime}[h]=g^{\prime\prime}(v)[h,f(v)]+g^{\prime}(v)[f^{\prime}(v)[h]]$
where indicates derivative w.r.t. $v$. It immediately follows that
$\partial_{s}\mathcal{G}^{1}\eta_{s}-f_{\textrm{W}}(\mathcal{G}^{1})(\eta_{s})$ $=-[f_{\textrm{W}},[f_{\textrm{W}},f_{\textrm{FR}}]](S_{\textrm{W}}(s,u_{\gamma}))=-\mathcal{G}^{2}\eta_{s},$
and thus
$f^{(1)}(s)$ $=-e^{(\gamma-s)f_{\textrm{W}}}\mathcal{G}^{2}\eta_{s}$
using the fact that $f_{\textrm{W}}$ and $e^{tf_{\textrm{W}}}$ commute for any $t>0$.
For $f^{(2)}$ we have
$f^{(2)}(s)$ $=f_{\textrm{W}}e^{(\gamma-s)f_{W}}\mathcal{G}^{2}\eta_{s}-e^{(\gamma-s)f_{W}}\partial_{s}\mathcal{G}^{2}\eta_{s}.$
Recalling that
$\mathcal{G}^{2}v=[f_{\textrm{W}},[f_{\textrm{W}},f_{\textrm{FR}}]](v)=f_{\textrm{W}}(\mathcal{G}^{1}v)-(\mathcal{G}^{1})^{\prime}(v)[f_{\textrm{W}}(v)],$
we find
$\partial_{s}f_{\textrm{W}}(\mathcal{G}^{1}\eta_{s})$ $=f_{\textrm{W}}(\mathcal{G}^{1}\eta_{s})[\partial_{s}\mathcal{G}^{1}\eta_{s}]=f_{\textrm{W}}(\mathcal{G}^{1}\eta_{s})\left([f_{\textrm{W}},f_{\textrm{FR}}]^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]\right)$
$\partial_{s}(\mathcal{G}^{1})^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]$ $=(\mathcal{G}^{1})^{\prime\prime}(\eta_{s})(f_{\textrm{W}}(\eta_{s}),f_{\textrm{W}}(\eta_{s}))+(\mathcal{G}^{1})^{\prime}(\eta_{s})[f_{\textrm{W}}^{2}(\eta_{s})].$
Therefore
$\partial_{s}\mathcal{G}^{2}\eta_{s}$ $=f_{\textrm{W}}(\mathcal{G}^{1}\eta_{s})\left([f_{\textrm{W}},f_{\textrm{FR}}]^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]\right)-(\mathcal{G}^{1})^{\prime\prime}(\eta_{s})(f_{\textrm{W}}(\eta_{s}),f_{\textrm{W}}(\eta_{s}))+(\mathcal{G}^{1})^{\prime}(\eta_{s})[f_{\textrm{W}}^{2}(\eta_{s})]$
$=f_{\textrm{W}}(\mathcal{G}^{1}\eta_{s})\left([f_{\textrm{W}},f_{\textrm{FR}}]^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]\right)-\left((\mathcal{G}^{1})^{\prime}(\eta_{s})f_{\textrm{W}}(\eta_{s})\right)^{\prime}[f_{\textrm{W}}(\eta_{s})]$
$=(\mathcal{G}^{2})^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]$
using the chain rule for operators.
Following the same steps as before we have $\partial_{s}\mathcal{G}^{2}\eta_{s}-f_{\textrm{W}}(\mathcal{G}^{2}\eta_{s})=-\mathcal{G}^{3}\eta_{s}$ and thus $f^{(2)}(s)=e^{(\gamma-s)f_{W}}\mathcal{G}^{3}\eta_{s}$. A recursive argument then gives
$f^{(k)}(s)$ $=(-1)^{k}e^{(\gamma-s)f_{W}}\mathcal{G}^{k+1}\eta_{s},$
from which follows
$e^{(\gamma-\tau)f_{\textrm{W}}}\mathcal{G}^{1}\eta_{\tau}=\mathcal{G}^{1}\eta_{\gamma}+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}(\tau-\gamma)^{k}\mathcal{G}^{k+1}\eta_{\gamma}.$
Integrating both sides w.r.t. $\tau$
$\int_{0}^{\gamma}e^{(\gamma-\tau)f_{\textrm{W}}}\mathcal{G}^{1}\eta_{\tau}d\tau$ $=\int_{0}^{\gamma}\mathcal{G}^{1}\eta_{\gamma}d\tau+\sum_{k=1}^{\infty}\frac{\mathcal{G}^{k+1}\eta_{\gamma}}{k!}\int_{0}^{\gamma}(\gamma-\tau)^{k}d\tau$
$=\gamma\mathcal{G}^{1}\eta_{\gamma}+\sum_{k=1}^{\infty}\frac{\mathcal{G}^{k+1}\eta_{\gamma}}{k!}\frac{\gamma^{k+1}}{k+1}$
$=\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}^{k}(\eta_{\gamma}).$
Lemma A.2.
Let Assumption 1 hold. Then the $k$th commutator $\mathcal{G}^{k}\eta_{\gamma}:=[f_{\textrm{W}},\mathcal{G}^{k-1}\eta_{\gamma}]$ for $k=1,2,3,\dots$ with $\mathcal{G}^{0}:=f_{\textrm{FR}}$ and $\mathcal{G}^{1}$ as defined in (2.17) has a Fisher-Rao structure,
$\mathcal{G}^{k}\eta_{\gamma}=\eta_{\gamma}\left(g_{k}(\eta_{\gamma})-\mathbb{E}_{\eta_{\gamma}}[g_{k}]\right)$
where
$g_{k}(\eta_{\gamma})$ $=-g_{k-1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})+\nabla\log\frac{\eta_{\gamma}}{\pi}\cdot\nabla g_{k-1}(\eta_{\gamma})+\frac{1}{\eta_{\gamma}}\nabla\cdot(\eta_{\gamma}\nabla g_{k-1}(\eta_{\gamma})),\quad\text{for}\enskip k=1,2,3,\dots$
$g_{0}(\eta_{\gamma})$ $=-\log\frac{\eta_{\gamma}}{\pi}$
and $g_{k-1}^{\prime}f_{\textrm{W}}(\eta)$ denotes the Fréchet derivative of $g_{k-1}$ in the direction $f_{\textrm{W}}(\eta)$.
Proof.
Recall $g_{0}=-\log\frac{\eta_{\gamma}}{\pi}$, where we drop the dependence on $\eta_{\gamma}$ to simplify notation, so that
$f_{\textrm{FR}}(\eta_{\gamma})=\eta_{\gamma}(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])$
We will proceed inductively, starting with $\mathcal{G}^{1}$,
$[f_{\textrm{W}},f_{\textrm{FR}}]\eta_{\gamma}$ $=f_{\textrm{W}}^{\prime}f_{\textrm{FR}}(\eta_{\gamma})-f_{\textrm{FR}}^{\prime}f_{\textrm{W}}(\eta_{\gamma}).$ (A.3)
Since $f_{\textrm{W}}$ is a bounded linear operator, its Fréchet derivative is itself and
$f_{\textrm{W}}^{\prime}f_{\textrm{FR}}(\eta_{\gamma})$ $=\nabla\cdot\left(\nabla\eta_{\gamma}(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])\right)-\nabla\cdot\left(\eta_{\gamma}(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])\nabla\log\pi\right)$ (A.4)
$=\nabla\cdot(\nabla\eta_{\gamma}g_{0})-\nabla\cdot(\eta_{\gamma}g_{0}\nabla\log\pi)-\mathbb{E}_{\eta_{\gamma}}[g_{0}]f_{\textrm{W}}(\eta_{\gamma})$
$=2\nabla g_{0}\cdot\nabla\eta_{\gamma}+\eta_{\gamma}\Delta g_{0}-\mu\nabla\log\pi\cdot\nabla g_{0}+(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])f_{\textrm{W}}(\eta_{\gamma})$
$=\nabla g_{0}\cdot\nabla\eta_{\gamma}+\eta_{\gamma}\Delta g_{0}+\left(\nabla\eta_{\gamma}-\eta_{\gamma}\nabla\log\pi\right)\cdot\nabla g_{0}+(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])f_{\textrm{W}}(\eta_{\gamma})$
$=\nabla\cdot(\eta_{\gamma}\nabla g_{0})+\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{0}+(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])f_{\textrm{W}}(\eta_{\gamma}).$
Notice that this expression holds for any suitably differentiable scalar valued function $g$ in place of $g_{0}$.
The Fréchet derivative of $f_{\textrm{FR}}$ can be obtained by computing its directional derivative
$\frac{d}{d\epsilon}\left.f_{\textrm{FR}}(\eta_{\gamma}+\epsilon h)\right|_{\epsilon=0}$ $=\frac{d}{d\epsilon}\left.-(\eta_{\gamma}+\epsilon h)\log\frac{\eta_{\gamma}+\epsilon h}{\pi}+(\eta_{\gamma}+\epsilon h)\int\log\frac{\eta_{\gamma}+\epsilon h}{\pi}(\eta_{\gamma}+\epsilon h)\right|_{\epsilon=0}$ (A.5)
$=\left(\log\frac{\pi}{\eta_{\gamma}}-1\right)h-h\int\log\frac{\pi}{\eta_{\gamma}}\eta_{\gamma}\,+\eta_{\gamma}\int h\left(1-\log\frac{\pi}{\eta_{\gamma}}\right)$
$=h\left(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}]\right)+\eta_{\gamma}\left(g_{0}^{\prime}h-\mathbb{E}_{\eta_{\gamma}}[g_{0}^{\prime}h]\right)-\eta_{\gamma}\mathbb{E}_{h}[g_{0}].$
So then since $g_{0}^{\prime}h=-\frac{h}{\eta_{\gamma}}$,
$f_{\textrm{FR}}^{\prime}f_{\textrm{W}}(\eta_{\gamma})$ $=f_{\textrm{W}}(\eta_{\gamma})\left(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}]\right)-\left(f_{\textrm{W}}(\eta_{\gamma})-\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[1]\right)-\eta_{\gamma}\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[g_{0}]$
$=f_{\textrm{W}}(\eta_{\gamma})\left(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}]\right)-\nabla\cdot\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)-\eta_{\gamma}\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[g_{0}]$
since $\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[1]=0$. Combining yields
$[f_{\textrm{W}},f_{\textrm{FR}}]\eta_{\gamma}$ $=(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])f_{\textrm{W}}(\eta_{\gamma})+\nabla\cdot(\eta_{\gamma}\nabla g_{0})+\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{0}$
$-f_{\textrm{W}}(\eta_{\gamma})\left(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}]\right)+\nabla\cdot\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)+\eta_{\gamma}\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[g_{0}]$
$=f_{\textrm{W}}(\eta_{\gamma})+\eta_{\gamma}\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[g_{0}]+\nabla\cdot(\eta_{\gamma}\nabla g_{0})+\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{0}$
$=-\eta_{\gamma}\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}-\eta_{\gamma}\int\log\frac{\eta_{\gamma}}{\pi}\nabla\cdot\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)$
$=-\eta_{\gamma}\left(\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}-\mathbb{E}_{\eta_{\gamma}}\left[\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}\right]\right)$ (A.6)
using the fact that $\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[1]=0$ in the second line and integration by parts in the last line. This yields the claim of the proof for $k=1$, since by the provided definition of $g_{k}$,
$g_{1}$ $=-g_{0}^{\prime}f_{\textrm{W}}(\eta_{\gamma})-\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}-\frac{1}{\eta_{\gamma}}\nabla\cdot\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)$
$=\frac{f_{\textrm{W}}(\eta_{\gamma})}{\eta_{\gamma}}-\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}-\frac{1}{\eta_{\gamma}}\nabla\cdot\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)$
$=-\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}.$
Since $\mathcal{G}^{1}$ has a FR structure, it is immediately clear that $\mathcal{G}^{2}$ will also, as
$[f_{\textrm{W}},[f_{\textrm{W}},f_{\textrm{FR}}]]=f_{\textrm{W}}^{\prime}[f_{\textrm{W}},f_{\textrm{FR}}](\eta_{\gamma})-[f_{\textrm{W}},f_{\textrm{FR}}]^{\prime}f_{\textrm{W}}(\eta_{\gamma}),$
which has the same structure as (A.3) with $g_{0}$ now replaced by $g_{1}$. Nevertheless, we will calculate it to show this is the case. The Fréchet derivative of $f_{\textrm{W}},f_{\textrm{FR}}]$ is given by
$\left.\frac{d}{d\epsilon}[f_{\textrm{W}},f_{\textrm{FR}}](\eta_{\gamma}+\epsilon h)\right|_{\epsilon=0}$ $=\left.\frac{d}{d\epsilon}-(\eta_{\gamma}+\epsilon h)\left(\left|\nabla\log\frac{\eta_{\gamma}+\epsilon h}{\pi}\right|^{2}-\int\left|\nabla\log\frac{\eta_{\gamma}+\epsilon h}{\pi}\right|^{2}(\eta_{\gamma}+\epsilon h)\right)\right|_{\epsilon=0}$
$=h(g_{1}-\mathbb{E}_{\eta_{\gamma}}[g_{1}])-\eta_{\gamma}\left(2\nabla\left(\frac{h}{\eta_{\gamma}}\right)\cdot\nabla\log\frac{\eta_{\gamma}}{\pi}-\int 2\nabla\left(\frac{h}{\eta_{\gamma}}\right)\cdot\nabla\log\frac{\eta_{\gamma}}{\pi}\eta_{\gamma}\right)-\eta_{\gamma}\int g_{1}h$
$=h(g_{1}-\mathbb{E}_{\eta_{\gamma}}[g_{1}])+\eta_{\gamma}(g_{1}^{\prime}h-\mathbb{E}_{\eta_{\gamma}}[g_{1}^{\prime}h])-\eta_{\gamma}\mathbb{E}_{h}[g_{1}]$
where $g_{1}^{\prime}h:=-2\nabla\left(\frac{h}{\eta_{\gamma}}\right)\cdot\nabla\log\frac{\eta_{\gamma}}{\pi}$ is the Fréchet derivative of $g_{1}$ in the direction $h$. Notice this is the same structure as (A.5) but with $g_{0}$ replaced by $g_{1}$. Similarly, we can directly use (A.4) with $g_{1}$ in place of $g_{0}$ for the term $f_{\textrm{W}}^{\prime}[f_{\textrm{W}},f_{\textrm{FR}}](\eta_{\gamma})$ due to the Fisher-Rao structure of $f_{\textrm{W}},f_{\textrm{FR}}]$. This yields
$\mathcal{G}^{2}\eta_{\gamma}$ $=f_{\textrm{W}}(\eta_{\gamma})(g_{1}-\mathbb{E}_{\eta_{\gamma}}[g_{1}])+\nabla\cdot(\eta_{\gamma}\nabla g_{1})+\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{1}$
$-\left(f_{\textrm{W}}(\eta_{\gamma})(g_{1}-\mathbb{E}_{\eta_{\gamma}}[g_{1}])+\eta_{\gamma}(g_{1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})-\mathbb{E}_{\eta_{\gamma}}[g_{1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})])-\eta_{\gamma}\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[g_{1}]\right)$
$=-\eta_{\gamma}(g_{1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})-\mathbb{E}_{\eta_{\gamma}}[g_{1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})])+\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{1}+\eta_{\gamma}\mathbb{E}_{f_{W}(\eta_{\gamma})}[g_{1}]+\nabla\cdot(\eta_{\gamma}\nabla g_{1})$
Notice that
$\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{1}+\eta_{\gamma}\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[g_{1}]=\eta_{\gamma}\left(\nabla\log\frac{\eta_{\gamma}}{\pi}\cdot\nabla g_{1}+\mathbb{E}_{\eta_{\gamma}}\left[-\nabla\log\frac{\eta_{\gamma}}{\pi}\cdot\nabla g_{1}\right]\right)$
using integration by parts for the expectation term. For the last term, we can re-write as
$\nabla\cdot(\eta_{\gamma}\nabla g_{1})=\eta_{\gamma}\left(\frac{1}{\eta_{\gamma}}\nabla\cdot(\eta_{\gamma}\nabla g_{1})-\mathbb{E}_{\eta_{\gamma}}\left[\frac{1}{\eta_{\gamma}}\nabla\cdot(\eta_{\gamma}\nabla g_{1})\right]\right)$
since, by the divergence theorem and the tail decay of $\eta_{\gamma},g_{1}$ guaranteed by Assumption 1, $\int\nabla\cdot(\eta_{\gamma}\nabla g_{1})=0$. Combining the above results yields
$\mathcal{G}^{2}\eta_{\gamma}=\eta_{\gamma}(g_{2}-\mathbb{E}_{\eta_{\gamma}}[g_{2}])$
with
$g_{2}=-g_{1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})+\nabla\log\frac{\eta_{\gamma}}{\pi}\cdot\nabla g_{1}+\frac{1}{\eta_{\gamma}}\nabla\cdot(\eta_{\gamma}\nabla g_{1})$
which confirms the claim of the lemma for $k=2$. By induction, the claim holds for all positive integer $k$.
Appendix B Splitting for multivariate Gaussian distributions
We recap here useful notation for this section:
  • $\mu_{0}(x)=\mathcal{N}(x;m_{0},C_{0})$ denotes the initial distribution and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$ the target distribution.
  • We denote by $\nu_{\gamma}$ the solution of the W-FR PDE in Proposition 2.2 for one time step $\gamma>0$.
  • We denote by $\eta_{\gamma}$ the solution of the FR-W PDE in Proposition 2.3 for one time step $\gamma>0$.
  • In the Gaussian case, $b_{\gamma},Q_{\gamma}$ denote mean and variance of $\nu_{\gamma}$.
  • In the Gaussian case, $a_{\gamma},P_{\gamma}$ denote mean and variance of $\eta_{\gamma}$.
B.1 Moment ODEs and analytic solutions
We collect here the derivation and exact solutions of the moment ODEs for WFR and sequential splitting. Before stating our results we recall the following properties of matrix exponentials: for $X,Y$ matrices we have
$e^{X}e^{Y}$ $=e^{X+Y}\quad\text{if}\quad XY=YX$
$e^{YXY^{-1}}$ $=Ye^{X}Y^{-1}\quad Y\enskip\text{invertible}$
$Xe^{X}$ $=e^{X}X.$
We first refine the result of [liero2025evolution, Remark 2.3] and obtain an explicit expression for the covariance matrix of the WFR flow rather than its inverse. The specific form derived below will be particularly useful to obtain the convergence rate of $C_{t}$ to $C_{\pi}$ in Lemma 3.1. Additionally, we obtain explicit solutions for the mean, which are not detailed in [liero2025evolution] and seem to be missing from the literature.
Throughout this section, we focus on PDEs of the form
$\partial_{t}\mu_{t}$ $=f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t})+\mu_{t}\left(g_{M}(\mu_{t})-\mathbb{E}_{\mu_{t}}[g_{M}]\right)$ (B.1)
$g_{M}(\mu_{t})$ $=\nabla g(\mu_{t})^{\top}M_{t}\nabla g(\mu_{t})$ (B.2)
initialised at $\mu_{0}=\mathcal{N}(x;m_{0},C_{0})$ with $g$ as defined in (2.4), $\pi=\mathcal{N}(x;m_{\pi},C_{\pi})$ and $M_{t}\in\mathbb{R}^{d\times d}$ a potentially time dependent symmetric matrix independent of $x,\mu_{t}$. The ODEs for the exact WFR given in [us, Appendix D] are obtained for $M_{t}\equiv 0$. Assume further that $C_{0}$ and $C_{\pi}$ are invertible (non-degenerate) covariance matrices. It holds straightforwardly that $\mu_{t}=\mathcal{N}(x;m_{t},C_{t})$ and the following lemmas characterise the evolution of $m_{t},C_{t}$ with time.
Lemma B.1.
Moment ODEs, Multivariate Gaussian. Consider (B.1) with the aforementioned conditions. Then $m_{t},C_{t}$ satisfy for all $t>0$,
$\frac{dC_{t}}{dt}$ $=-C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})C_{t}-\left(I-\tfrac{1}{2}C_{\pi}+2M_{t}\right)C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}(I-\tfrac{1}{2}C_{\pi}+2M_{t})+2(I+M_{t})$ (B.3)
$\frac{dm_{t}}{dt}$ $=-(C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+(I+2M_{t})C_{\pi}^{-1})(m_{t}-m_{\pi})$ (B.4)
Proof.
By straightforward integration and assuming interchange of derivatives and integrals,
$\frac{dC_{t}}{dt}=\int(x-m_{t})(x-m_{t})^{\top}\partial_{t}\mu_{t}(x)dx$ $=\int(x-m_{t})(x-m_{t})^{\top}(f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}))dx$ (B.5)
$+\int(x-m_{t})(x-m_{t})^{\top}(g_{M}(\mu_{t})-\mathbb{E}_{\mu_{t}}[g_{M}])\mu_{t}(x)dx.$
The first term on the rhs is obtained directly from the covariance ODE for the exact WFR, see [us, Appendix D] for a derivation,
$\int(x-m_{t})(x-m_{t})^{\top}(f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}))dx=-C_{t}C_{\pi}^{-1}C_{t}+C_{t}-C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}+2I.$ (B.6)
The remaining term is straightforwardly evaluated using that for $\mu_{t}(x)=\mathcal{N}(x;m_{t},C_{t})$,
$g_{M}(\mu_{t})=(x-m_{t}+m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(x-m_{t}+m_{t}-\tilde{m}_{t})$ (B.7)
where $\tilde{C}_{t}^{-1}=C_{\pi}^{-1}-C_{t}^{-1}$ and $\tilde{m}_{t}=\tilde{C}_{t}(C_{\pi}^{-1}m_{\pi}-C_{t}^{-1}m_{t})$. Then,
$\int(x-m_{t})(x-m_{t})^{\top}g_{M}(\mu_{t})\mu_{t}(x)dx=\int(x-m_{t})(x-m_{t})^{\top}(x-m_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(x-m_{t})\mu_{t}(x)dx$
$+2\int(x-m_{t})(x-m_{t})^{\top}(x-m_{t})^{\top}\mu_{t}(x)dx\,\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$
$+C_{t}(m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$
$=2C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}C_{t}+C_{t}Tr[C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}]+C_{t}(m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$
$=2(C_{t}C_{\pi}^{-1}M_{t}C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}M_{t}-M_{t}C_{\pi}^{-1}C_{t}+M_{t})+C_{t}Tr[C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}]$
$+C_{t}(m_{t}-\tilde{m}_{\gamma})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$ (B.8)
where in the second last line we have used that for a random variable $X\in\mathbb{R}^{d\times 1}$, $X\sim N(0,\Sigma)$ and any constant matrix $A$, $\mathbb{E}[XX^{\top}X^{\top}AX]=\mathbb{E}[XX^{T}AXX^{T}]=\Sigma(A+A^{\top})\Sigma+\Sigma Tr[\Sigma A]$. Also,
$\mathbb{E}_{\mu_{t}}[g_{M}]\int(x-m_{t})(x-m_{t})^{\top}\mu_{t}(x)dx=C_{t}\left(Tr[C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}]+(m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})\right)$ (B.9)
Substituting (B.8) and (B.9) into (B.5) and combining with (B.6) yields the covariance ODE (B.3).
The mean ODE is obtained in a similar way
$\frac{dm_{t}}{dt}=\int x\partial_{t}\mu_{t}(x)dx$ $=\int x(f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}))dx+\int x(g_{M}(\mu_{t})-\mathbb{E}_{\mu_{t}}[g_{M}])\mu_{t}(x)dx,$
where the first term on the rhs is obtained directly from the mean ODE for the exact WFR, see [us, Appendix D] for derivation,
$\int x(f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}))dx=-(C_{t}+I)C_{\pi}^{-1}(m_{t}-m_{\pi})$ (B.10)
and for the remaining term, using (B.7),
$\int x(g_{M}(\mu_{t})-\mathbb{E}_{\mu_{t}}[g_{M}])\mu_{t}(x)dx$ $=\int(x-m_{t})(x-m_{t}+m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(x-m_{t}+m_{t}-\tilde{m}_{t})\mu_{t}(x)dx$
$=2C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$
$=2C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{C}_{t}C_{\pi}^{-1}m_{\pi}+\tilde{C}_{t}C_{t}^{-1}m_{t})$
$=2C_{t}\tilde{C}_{t}^{-1}M_{t}C_{\pi}^{-1}m_{t}-2C_{t}\tilde{C}_{t}^{-1}M_{t}C_{\pi}^{-1}m_{\pi}$
$=2(C_{t}C_{\pi}^{-1}-I)M_{t}C_{\pi}^{-1}(m_{t}-m_{\pi}).$
Combining with (B.10) yields (B.4).
Lemma B.2.
Analytic Solutions, Multivariate Gaussian. Given the conditions of Lemma B.1 and $\Gamma:=C_{\pi}^{-1}+\tfrac{1}{2}I$. Then the mean and covariance ODEs (B.4) and (B.3) with initial conditions $m_{0},C_{0}$ have analytic solution for all $t>0$ given by
$C_{t}$ $=C_{\pi}+e^{-\Gamma t}E_{0}(I-(W_{t}^{-1}+E_{0})^{-1}E_{0})e^{-\Gamma t}$ (B.11)
$m_{t}$ $=m_{\pi}+(C_{t}-C_{\pi})e^{tC_{\pi}^{-1}}(C_{0}-C_{\pi})^{-1}(m_{0}-m_{\pi}).$ (B.12)
where $W_{t}:=(2I+C_{\pi})^{-1}-Z_{t}^{M}-e^{-\Gamma t}(2I+C_{\pi})^{-1}e^{-\Gamma t}$, $Z_{t}^{M}:=2\int_{0}^{t}e^{-\Gamma s}M_{s}e^{-\Gamma s}ds$ with $Z_{t}^{M}<\infty$ for all $t>0$ and $E_{0}:=C_{0}-C_{\pi}$. Notice in particular that $m_{t}$ is independent of $M_{t}$.
An alternative presentation for the covariance for the case $C_{0}\neq C_{\pi}$ is given by
$C_{t}=C_{\pi}+(e^{\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}]e^{\Gamma t}-(2I+C_{\pi})^{-1})^{-1}.$ (B.13)
Proof.
Beginning with the mean ODE, consider the transformation
$z_{t}=(C_{t}-C_{\pi})^{-1}(m_{t}-m_{\pi})$ (B.14)
and differentiating with respect to $t$ yields (using the shorthand notation $\frac{dC_{t}}{dt}=\dot{C}_{t}$,
$\dot{z_{t}}$ $=-(C_{t}-C_{\pi})^{-1}\dot{C}_{t}(C_{t}-C_{\pi})^{-1}(m_{t}-m_{\pi})+(C_{t}-C_{\pi})^{-1}\dot{m}_{t}$
$=-(C_{t}-C_{\pi})^{-1}\dot{C}_{t}z_{t}-(C_{t}-C_{t})^{-1}(C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+(I+2M_{t})C_{\pi}^{-1})(C_{t}-C_{\pi})z_{t}$
using (B.4) and $(C_{t}-C_{\pi})z_{t}=m_{t}-m_{\pi}$. Rearranging (B.3) as
$\dot{C}_{t}$ $=(2C_{t}C_{\pi}^{-1}M_{t}C_{\pi}^{-1}-2M_{t}C_{\pi}^{-1}-C_{t}C_{\pi}^{-1}-C_{\pi}^{-1})(C_{t}-C_{\pi})-(C_{t}-C_{\pi})C_{\pi}^{-1}$
$=(C_{t}C_{\pi}^{-1}(2M_{t}C_{\pi}^{-1}-I)-(I+2M_{t})C_{\pi}^{-1})(C_{t}-C_{\pi})-(C_{t}-C_{\pi})C_{\pi}^{-1}$
and substituting into the ODE for $z_{t}$ yields
$\dot{z_{t}}$ $=-(C_{t}-C_{\pi})^{-1}[C_{t}C_{\pi}^{-1}(2M_{t}C_{\pi}^{-1}-I)-(I+2M_{t})C_{\pi}^{-1}](C_{t}-C_{\pi})z_{t}+(C_{t}-C_{\pi})^{-1}(C_{t}-C_{\pi})C_{\pi}^{-1}z_{t}$
$-(C_{t}-C_{t})^{-1}(C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+(I+2M_{t})C_{\pi}^{-1})(C_{t}-C_{\pi})z_{t}$
$=C_{\pi}^{-1}z_{t}$
which has analytic solution $z_{t}=e^{tC_{\pi}^{-1}}z_{0}$. Inserting (B.14) and rearranging yields (B.12).
For the covariance ODE, we first consider the so-called natural transformation as in [liero2025evolution] to obtain an analytic solution for $C_{t}^{-1}$. Using the transformation $A_{t}=C_{t}^{-1}$ and differentiating with respect to time yields
$\dot{A}_{t}$ $=-A_{t}\dot{C}_{t}A_{t}$
$=-C_{t}^{-1}\left(-C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})C_{t}-\left(I-\tfrac{1}{2}C_{\pi}+2M_{t}\right)C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}(I-\tfrac{1}{2}C_{\pi}+2M_{t})+2(I+M_{t})\right)C_{t}^{-1}$
$=C_{\pi}^{-1}-2C_{\pi}^{-1}M_{t}C_{\pi}^{-1}+C_{t}^{-1}\left(I+2M_{t}\right)C_{\pi}^{-1}+C_{\pi}^{-1}(I+2M_{t})C_{t}^{-1}-C_{t}^{-1}-2C_{t}^{-1}(I+M_{t})C_{t}^{-1}$
$=C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+A_{t}\left(I+2M_{t}\right)C_{\pi}^{-1}+C_{\pi}^{-1}(I+2M_{t})A_{t}-A_{t}-2A_{t}(I+M_{t})A_{t}$
Let $\Sigma_{t}:=A_{t}-C_{\pi}^{-1}$ and $\Gamma:=C_{\pi}^{-1}+\frac{1}{2}I$ and $B_{t}:=e^{\Gamma t}\Sigma_{t}e^{\Gamma t}$, then
$\dot{B}_{t}=e^{\Gamma t}(\Gamma\Sigma_{t}+\dot{\Sigma}_{t}+\Sigma_{t}\Gamma)e^{\Gamma t}$ (B.15)
and
$\Gamma\Sigma_{t}+\dot{\Sigma}_{t}+\Sigma_{t}\Gamma$ $=\left(C_{\pi}^{-1}+\tfrac{1}{2}I\right)(A_{t}-C_{\pi}^{-1})+(A_{t}-C_{\pi}^{-1})\left(C_{\pi}^{-1}+\tfrac{1}{2}I\right)$
$+C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+A_{t}\left(I+2M_{t}\right)C_{\pi}^{-1}+C_{\pi}^{-1}(I+2M_{t})A_{t}-A_{t}-2A_{t}(I+M_{t})A_{t}$
$=C_{\pi}^{-1}A_{t}+A_{t}C_{\pi}^{-1}-2C_{\pi}^{-2}-C_{\pi}^{-1}+A_{t}+A_{t}\left(I+2M_{t}\right)C_{\pi}^{-1}+C_{\pi}^{-1}(I+2M_{t})A_{t}$
$+C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})-A_{t}-2A_{t}(I+M_{t})A_{t}$
$=-2\Sigma_{t}(I+M_{t})\Sigma_{t}$
and inserting into (B.15) yields
$\dot{B}_{t}$ $=-2e^{\Gamma t}\Sigma_{t}(I+M_{t})\Sigma_{t}e^{\Gamma t}=-2B_{t}e^{-\Gamma t}(I+M_{t})e^{-\Gamma t}B_{t}$
Consider a further transformation, $D_{t}=B_{t}^{-1}$, then
$\dot{D}_{t}=-D_{t}\dot{B}_{t}D_{t}=2e^{-\Gamma t}(I+M_{t})e^{-\Gamma t}$ (B.16)
whose analytic solution has the representation
$D_{t}=D_{0}+2\int_{0}^{t}e^{-\Gamma s}(I+M_{s})e^{-\Gamma s}ds=(C_{0}^{-1}-C_{\pi}^{-1})^{-1}+Z_{t}$
with $Z_{t}:=2\int_{0}^{t}e^{-\Gamma s}(I+M_{s})e^{-\Gamma s}ds$ and since $D_{0}=(C_{0}^{-1}-C_{\pi}^{-1})^{-1}$. Finally,
$C_{t}^{-1}=C_{\pi}^{-1}+\Sigma_{t}=C_{\pi}^{-1}+e^{-\Gamma t}D_{t}^{-1}e^{-\Gamma t}=C_{\pi}^{-1}+e^{-\Gamma t}((C_{0}^{-1}-C_{\pi}^{-1})^{-1}+Z_{t})^{-1}e^{-\Gamma t}.$ (B.17)
Note that for the special case $M_{t}=0$, $Z_{t}=2\int_{0}^{t}e^{-2\Gamma s}ds=\Gamma^{-1}(I-e^{-2\Gamma t})$ so that the above coincides with the expression for $C_{t}^{-1}$ in [liero2025evolution, Remark 2.3] with $\alpha=\beta=1$ there.
This representation however is not ideal as it forces the condition $C_{0}^{-1}\neq C_{\pi}^{-1}$ to ensure invertibility of their difference. Note also that $C_{0}=C_{\pi}$ is a valid initial condition, wherein $\frac{dC_{t}}{dt}=0$ for all $t>0$ and (B.4) remains well-posed. We will consider an alternative representation, by first writing (B.17) in terms of $C_{t}$ using the Woodbury matrix identity,
$(A+UWV)^{-1}=A^{-1}-A^{-1}U(W^{-1}+VA^{-1}U)^{-1}VA^{-1},$ (B.18)
with $A=C_{\pi}^{-1},W=(D_{0}+Z_{t})^{-1},\,U=V=e^{-\Gamma t}$ yields
$C_{t}$ $=C_{\pi}-C_{\pi}e^{-\Gamma t}(D_{0}+Z_{t}+e^{-2\Gamma t}C_{\pi})^{-1}e^{-\Gamma t}C_{\pi}$
$=C_{\pi}+(R_{t}^{-1}X_{t}R_{t}^{-1})^{-1}$
where $R_{t}:=C_{\pi}e^{-\Gamma t}$ and $X_{t}:=(C_{\pi}^{-1}-C_{0}^{-1})^{-1}-Z_{t}-e^{-\Gamma t}C_{\pi}e^{-\Gamma t}$. Then let $Z_{t}=Z_{t}^{M}+\Gamma^{-1}(I-e^{-2\Gamma t})$ with $Z_{t}^{M}:=2\int_{0}^{t}e^{-\Gamma s}M_{s}e^{-\Gamma s}ds$. Since $(C_{\pi}^{-1}-C_{0}^{-1})^{-1}=C_{\pi}(C_{0}-C_{\pi})^{-1}C_{0}$, we have
$R_{t}^{-1}X_{t}R_{t}^{-1}$ $=e^{\Gamma t}C_{\pi}^{-1}((C_{\pi}^{-1}-C_{0}^{-1})^{-1}-Z_{t}-e^{-\Gamma t}C_{\pi}e^{-\Gamma t})C_{\pi}^{-1}e^{\Gamma t}$
$=e^{\Gamma t}((C_{0}-C_{\pi})^{-1}C_{0}C_{\pi}^{-1}-C_{\pi}^{-1}(Z_{t}^{M}+\Gamma^{-1}(I-e^{-2\Gamma t}))C_{\pi}^{-1})e^{\Gamma t}-C_{\pi}^{-1}$
$=e^{\Gamma t}((C_{0}-C_{\pi})^{-1}C_{0}C_{\pi}^{-1}-C_{\pi}^{-2}\Gamma^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1})e^{\Gamma t}-C_{\pi}^{-1}+\Gamma^{-1}C_{\pi}^{-2}$ (B.19)
using that $C_{\pi},\,\Gamma$ and $e^{-\Gamma t}$ commute. We can further simplify using Woodbury identities, beginning with
$-C_{\pi}^{-1}+C_{\pi}^{-2}\Gamma^{-1}$ $=-C_{\pi}^{-1}+C_{\pi}^{-2}(C_{\pi}^{-1}+\tfrac{1}{2}I)^{-1}$
$=-C_{\pi}^{-1}+((C_{\pi}^{-1}+\tfrac{1}{2}I)C_{\pi}^{2})^{-1}$
$=-\frac{1}{2}C_{\pi}^{-1}C_{\pi}\left(I+\tfrac{1}{2}C_{\pi}\right)^{-1}$
$=-(2I+C_{\pi})^{-1}$ (B.20)
where in the second last line, we have used the special case of the Woodbury identity
$(I+A)^{-1}-I=-A(I+A)^{-1}$ (B.21)
with $A=\frac{1}{2}C_{\pi}$. Now consider, by similar reasoning,
$(C_{0}-C_{\pi})^{-1}C_{0}C_{\pi}^{-1}-\Gamma^{-1}C_{\pi}^{-2}$ $=(C_{0}-C_{\pi})^{-1}C_{0}C_{\pi}^{-1}-\left(I+\tfrac{1}{2}C_{\pi}\right)^{-1}C_{\pi}^{-1}$
$=\left((C_{0}-C_{\pi})^{-1}C_{0}-I+\tfrac{1}{2}C_{\pi}(I+\tfrac{1}{2}C_{\pi})^{-1}\right)C_{\pi}^{-1}$
$=\left((C_{0}-C_{\pi})^{-1}C_{0}-I\right)C_{\pi}^{-1}+(2I+C_{\pi})^{-1}$
$=(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}$ (B.22)
again using (B.21), $(I+\frac{1}{2}C_{\pi})^{-1}=-I+\frac{1}{2}C_{\pi}(I+\frac{1}{2}C_{\pi})^{-1}$ in the second and third line, and then using (B.18) in the fourth line with $A=-C_{\pi},U=C_{\pi},V=C_{0},W=(C_{\pi})^{-1}$ so that
$(C_{0}-C_{\pi})^{-1}=-C_{\pi}^{-1}-(C_{\pi}-C_{0})^{-1}C_{0}C_{\pi}^{-1}$ (B.23)
Substituting (B.20) and (B.22) into (B.19) yields
$R_{t}^{-1}X_{t}R_{t}^{-1}=e^{\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}]e^{\Gamma t}-(2I+C_{\pi})^{-1}$
which then yields the expression
$C_{t}=C_{\pi}+(e^{\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}]e^{\Gamma t}-(2I+C_{\pi})^{-1})^{-1}.$
This form will prove useful in the analysis of $\mathrm{KL}$ decay in Proposition 3.2. Finally, with another application of the Woodbury identity (B.18) with $A=(C_{0}-C_{\pi})^{-1},\,U=V=I,\,W=(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}-e^{-\Gamma t}(2I+C_{\pi})^{-1}e^{-\Gamma t}$, we obtain the representation
$C_{t}$ $=C_{\pi}+(e^{\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}-e^{-\Gamma t}(2I+C_{\pi})^{-1}e^{-\Gamma t}]e^{\Gamma t})^{-1}$
$=C_{\pi}+e^{-\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}-e^{-\Gamma t}(2I+C_{\pi})^{-1}e^{-\Gamma t}]^{-1}e^{-\Gamma t}$
$=C_{\pi}+e^{-\Gamma t}((C_{0}-C_{\pi})-(C_{0}-C_{\pi})(W^{-1}+(C_{0}-C_{\pi}))^{-1}(C_{0}-C_{\pi}))e^{-\Gamma t}$
as claimed.
B.2 Derivation of (2.20)
Using the representation (2.14), we will start by calculating the form of $g_{k}$ for $k=2,3,\dots$, and recall from Proposition 2.3 that
$\partial_{\gamma}\eta_{\gamma}=f_{\textrm{W}}(\eta_{\gamma})+f_{\textrm{FR}}(\eta_{\gamma})+\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}_{k}\eta_{\gamma}$
where
$\mathcal{G}_{k}\eta$ $=\eta\left(g_{k}(\eta)-\mathbb{E}_{\eta}[g_{k}]\right)$
$g_{k}(\eta)$ $=-g_{k-1}^{\prime}f_{W}(\eta)+\nabla\log\frac{\eta}{\pi}\cdot\nabla g_{k-1}(\eta)+\frac{1}{\eta}\nabla\cdot(\eta\nabla g_{k-1}(\eta))$
$=-g_{k-1}^{\prime}f_{W}(\eta)+(2\nabla g(\eta)+\nabla\log\pi)^{\top}\nabla g_{k-1}(\eta)+\Delta g_{k-1}(\eta),\enskip k=1,2,3,\dots$
$g_{0}(\eta)$ $=-\log\frac{\eta}{\pi}=-g(\eta)$
where $g$ is as defined in (2.4).
[…]
For $g_{1}$, we have that the Fréchet derivative term is given by $g_{0}^{\prime}f_{\textrm{W}}(\eta)=-\frac{f_{\textrm{W}}}{\eta}$ so that
$g_{1}=\frac{f_{\textrm{W}}(\eta)}{\eta}-|\nabla g_{0}|^{2}-\frac{f_{\textrm{W}}(\eta)}{\eta}=-|\nabla g|^{2}$
For $g_{2}$,
$\nabla g_{1}=-2H\nabla g\,;\quad\Delta g_{1}=-2Tr[H^{2}]$
and the Fréchet derivative term is given by
$g_{1}^{\prime}h$ $=-2\left(\nabla\frac{h}{\eta}\right)^{\top}\nabla g$
$g_{1}^{\prime}f_{\textrm{W}}$ $=-2\left(\nabla\frac{\nabla\cdot(\eta\nabla g)}{\eta}\right)^{\top}\nabla g$
$=-2\left(\nabla(\nabla\log\eta^{\top}\nabla g+\Delta g)\right)^{\top}\nabla g$
$=-2\left(\nabla(\nabla g^{\top}\nabla g+\nabla\log\pi^{\top}\nabla g)\right)^{\top}\nabla g$
$=-2\left(2H\nabla g+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}\nabla g$
$=-2\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}\nabla g$
since $\Delta g=Tr[H]$. Combining,
$g_{2}$ $\stackrel{{\scriptstyle+}}{{=}}2\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}\nabla g+(2\nabla g+\nabla\log\pi)^{\top}\nabla g_{1}$
$=2\nabla g^{\top}H_{\pi}\nabla g$
Similarly, for $g_{3}$,
$\nabla g_{2}=4HH_{\pi}\nabla g\,;\quad\Delta g_{2}=4Tr[HH_{\pi}H]$
using the identity $\nabla{\bf v}^{\top}A{\bf w}=\nabla{\bf v}^{\top}A{\bf w}+\nabla{\bf w}^{\top}A^{\top}{\bf v}$ for any ${\bf v}(x),{\bf w}(x):\mathbb{R}^{d\times 1}\rightarrow\mathbb{R}^{d\times 1}$ and $A\in\mathbb{R}^{d\times d}$ constant matrix independent of $x$ (see (93) in [Pedersen2012]). Then for the Fréchet derivative,
$g_{2}^{\prime}h$ $=4\left(\nabla\frac{h}{\eta}\right)^{\top}H_{\pi}\nabla g$
$g_{2}^{\prime}f_{\textrm{W}}$ $=4\left(\nabla\frac{h}{f_{\textrm{W}}}\right)^{\top}H_{\pi}\nabla g=4\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}\nabla g$
by the same reasoning as earlier. Combining,
$g_{3}$ $\stackrel{{\scriptstyle+}}{{=}}-4\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}\nabla g+(2\nabla g+\nabla\log\pi)^{\top}\nabla g_{2}$
$=-4\left(2H\nabla g+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}\nabla g+(2\nabla g+\nabla\log\pi)^{\top}4HH_{\pi}\nabla g$
$=\nabla g^{\top}(-8HH_{\pi}-4H_{\pi}^{2}+8HH_{\pi})\nabla g+\nabla\log\pi^{\top}(-4HH_{\pi}+4HH_{\pi})\nabla g$
$=-4\nabla g^{\top}H_{\pi}^{2}\nabla g$
and finally for $g_{4}$, by similar line of reasoning as earlier,
$\nabla g_{3}=-8HH_{\pi}^{2}\nabla g\,;\quad\Delta g_{3}=-8Tr[HH_{\pi}^{2}H]$
and the Fréchet derivative term is given by,
$g_{3}^{\prime}h$ $=-8\left(\nabla\frac{h}{\mu}\right)^{\top}H_{\pi}^{2}\nabla g$
$g_{3}^{\prime}f_{\textrm{W}}$ $=-8\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}^{2}\nabla g$
Combining,
$g_{4}$ $\stackrel{{\scriptstyle+}}{{=}}8\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}^{2}\nabla g+(2\nabla g+\nabla\log\pi)^{\top}\nabla g_{3}$
$=8\left(2H\nabla g+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}^{2}\nabla g-8(2\nabla g+\nabla\log\pi)^{\top}HH_{\pi}^{2}\nabla g$
$=8\nabla g^{\top}H_{\pi}^{3}\nabla g$
By induction, it holds that
$g_{k}$ $=-\nabla g^{\top}(-2H_{\pi})^{k-1}\nabla g=-\nabla g^{\top}(2C_{\pi}^{-1})^{k-1}\nabla g$
Finally, we have by the series representation of a matrix exponential that
$\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}_{k}\eta_{k}$ $=-\eta\nabla g^{\top}\left(\frac{C_{\pi}}{2}\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}(2C_{\pi}^{-1})^{k}\right)\nabla g+\eta\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\int\nabla g^{\top}(2C_{\pi}^{-1})^{k-1}\nabla g\eta dx$
$=-\eta\nabla g^{\top}\left(\frac{C_{\pi}}{2}\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}(2C_{\pi}^{-1})^{k}\right)\nabla g+\eta\int\nabla g^{\top}\left(\frac{C_{\pi}}{2}\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}(2C_{\pi}^{-1})^{k}\right)\nabla g\eta dx$
$=-\eta\left(g_{\text{FRW}}-\mathbb{E}_{\eta}[g_{\text{FRW}}]\right)$
where
$g_{\text{FRW}}:=\nabla g^{\top}\frac{C_{\pi}}{2}\left(e^{2\gamma C_{\pi}^{-1}}-I\right)\nabla g.$
B.3 Proof of Lemma 3.1
Due to Gaussianity, we have by a simple algebraic manipulation that
$\mathrm{KL}(\nu_{n}||\pi)$ $=\frac{1}{2}\left[-\log\frac{\det Q_{n\gamma}}{\det C_{\pi}}+(m_{n}^{\beta}-m_{\pi})^{\top}C_{\pi}^{-1}(m_{n}^{\beta}-m_{\pi})+Tr[C_{\pi}^{-1}Q_{n\gamma}]-d\right]$ (B.24)
$=\frac{1}{2}\left[-\log[\det(I+C_{\pi}^{-1}E_{n}^{\beta})]+(\varepsilon_{n}^{\beta})^{\top}C_{\pi}^{-1}\varepsilon_{n}^{\beta}+Tr[C_{\pi}^{-1}E_{n}^{\beta}]\right],$
where $E_{n}^{\beta}$ and $\varepsilon_{n}^{\beta}$ are as defined in (3.7) and (3.8) respectively, $\det(I+C_{\pi}^{-1}E_{n}^{\beta})>0$ and since $\det(A+B)=\det(A)\det(I+A^{-1}B)$ for square matrices $A,B$ with $A$ invertible. A similar expression holds for both $\mathrm{KL}(\mu_{n\gamma}||\pi)$ and $\mathrm{KL}(\eta_{n\gamma}||\pi)$ with $Q_{n\gamma}$ replaced by $C_{n\gamma}$ and $P_{n\gamma}$ respectively.
Applying the single step covariance equations derived in (3.3) recursively, one readily obtains the covariance recursion for the W-FR scheme,
$Q_{(n+1)\gamma}$ $=C_{\pi}+(e^{\Gamma\gamma}(Q_{n\gamma}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1})^{-1},$ (B.25)
with $\Gamma=C_{\pi}^{-1}+\frac{1}{2}I$. Likewise, one obtains for the FR-W scheme,
$P_{(n+1)\gamma}=C_{\pi}+\left(e^{\Gamma\gamma}(P_{n\gamma}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1}e^{2\gamma C_{\pi}^{-1}}\right)^{-1}$ (B.26)
Compare these with the recursion obtained from the solving the WFR PDE as derived in (3.1) over time interval $n\gamma,(n+1)\gamma]$, (N.B. in the below, we drop the $\gamma$ from the subscript where there is no ambiguity),
$C_{(n+1)\gamma}=C_{\pi}+(e^{\Gamma\gamma}[(C_{n\gamma}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}]e^{\Gamma\gamma}-(2I+C_{\pi})^{-1})^{-1}.$ (B.27)
Rearranging (B.25), (B.26) and (B.27) yields
$E_{n+1}^{\beta}$ $=(e^{\Gamma\gamma}(E_{n}^{\beta})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1})^{-1}$
$E_{n+1}^{\alpha}$ $=(e^{\Gamma\gamma}(E_{n}^{\alpha})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1}e^{2\gamma C_{\pi}^{-1}})^{-1}$
$E_{n+1}$ $=(e^{\Gamma\gamma}[E_{n}^{-1}+(2I+C_{\pi})^{-1}]e^{\Gamma\gamma}-(2I+C_{\pi})^{-1})^{-1}$
respectively, and by induction,
$E_{n}^{\beta}$ $=\left(e^{n\gamma\Gamma}E_{0}^{-1}e^{n\gamma\Gamma}+(e^{\gamma}-1)C_{\pi}^{-1}\sum_{j=0}^{n-1}e^{2j\gamma\Gamma}\right)^{-1}$
$=\left(e^{n\gamma\Gamma}E_{0}^{-1}e^{n\gamma\Gamma}+(1-e^{\gamma})C_{\pi}^{-1}(I-e^{2\gamma\Gamma})^{-1}(e^{2\gamma n\Gamma}-I)\right)^{-1}$
$=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+(1-e^{\gamma})(I-e^{2\gamma\Gamma})^{-1}C_{\pi}^{-1}(I-e^{-2\gamma n\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$ (B.28)
$E_{n}^{\alpha}$ $=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+(1-e^{\gamma})(e^{-2\gamma C_{\pi}^{-1}}-e^{\gamma})^{-1}C_{\pi}^{-1}(I-e^{-2\gamma n\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$ (B.29)
$E_{n}$ $=\left(e^{n\gamma\Gamma}E_{0}^{-1}e^{n\gamma\Gamma}+\frac{1}{2}\Gamma^{-1}C_{\pi}^{-1}(e^{2n\gamma\Gamma}-I)\right)^{-1}$
$=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+\frac{1}{2}\Gamma^{-1}C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$ (B.30)
since $\Gamma,e^{\Gamma},C_{\pi}$ all commute. In (B.28) and (B.29) we have used that $\sum_{j=0}^{n-1}T^{j}=(I-T)^{-1}(I-T^{n})$ for (IT)(I-T) invertible with $T=e^{2\gamma\Gamma}$.
The remaining term involving $\varepsilon_{n}$ is obtained by rearranging (3.3), (3.6) and (3.1) to obtain
$\varepsilon_{n}^{\beta}$ $=E_{n}^{\beta}e^{\gamma C_{\pi}^{-1}}(E_{n-1}^{\beta})^{-1}\varepsilon_{n-1}^{\beta}$
$\varepsilon_{n}^{\alpha}$ $=E_{n}^{\alpha}e^{\gamma C_{\pi}^{-1}}(E_{n-1}^{\alpha})^{-1}\varepsilon_{n-1}^{\alpha}$
$\varepsilon_{n}$ $=E_{n}e^{\gamma C_{\pi}^{-1}}E_{n-1}^{-1}\varepsilon_{n-1}$
respectively. By induction, one obtains that,
$(\varepsilon_{n}^{\beta})^{\top}C_{\pi}^{-1}\varepsilon_{n}^{\beta}=\varepsilon_{0}^{\top}E_{0}^{-1}e^{n\gamma C_{\pi}^{-1}}E_{n}^{\beta}C_{\pi}^{-1}E_{n}^{\beta}e^{n\gamma C_{\pi}^{-1}}E_{0}^{-1}\varepsilon_{0}$
and the same expression holds for $(\varepsilon_{n}^{\alpha})^{\top}C_{\pi}^{-1}\varepsilon_{n}^{\alpha}$ and $(\varepsilon_{n})^{\top}C_{\pi}^{-1}\varepsilon_{n}$ with $E_{n}^{\beta}$ replaced by $E_{n}^{\alpha}$ and $E_{n}$ respectively.
Combining all the derived expressions yields the result.
B.4 Proof of Proposition 3.2
(i). Begin with the case where $E_{0}$ is strictly positive definite. Recall $E_{n}=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+\Omega C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$. It trivially holds that $E_{n}\succ 0$ for all $n=1,2,3\dots$ since $AMA\succ 0$ if $A,M\succ 0$, which is true when letting $A=e^{-n\gamma\Gamma}$ and $M=E_{0}^{-1}+\Omega C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})$, where the latter is symmetric as $E_{0}^{-1}$ is the inverse of a symmetric matrix and $\Omega,C_{\pi},\Gamma$ are symmetric and they all commute with each other. The same reasoning holds for the claim for $E_{n}^{\beta}$ and $E_{n}^{\alpha}$.
(ii). For the second case, notice that when $E_{0}^{-1}+\Omega C_{\pi}^{-1}\prec 0$, then $E_{0}^{-1}+\Omega C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\prec 0$ and $(E_{0}^{-1}+\Omega C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma}))^{-1}\prec 0$ since $\Omega\succ 0$. The claim then holds since $AMA\prec 0$ for all positive definite $A,M$ with $A=e^{-n\gamma\Gamma}$ and $M=(E_{0}^{-1}+\Omega C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma}))^{-1}$. The same reasoning holds for $E_{n}^{\beta},E_{n}^{\alpha}$ with $\Omega$ replaced by $\Omega^{\beta}$ and $\Omega^{\alpha}$ respectively.
(iii). Note in the below, with a slight abuse of notation, we write $A_{n}$ to denote the value of the matrix $A$ at the $n$th iteration and $A_{ij}$ for the $i$th row and $j$th column element of the matrix $A$. We will demonstrate the reasoning for $\mathrm{KL}(\mu_{n\gamma}||\pi)$, noting that the same calculations hold for $\mathrm{KL}(\nu_{n}||\pi)$ and $\mathrm{KL}(\eta_{n}||\pi)$ due to Lemma 3.1 with $E_{n}$ replaced by $E_{n}^{\beta}$ and $E_{n}^{\alpha}$ respectively.
Let $A_{n}:=C_{\pi}^{-1}E_{n}$ and using Lemma 3.1 and (B.24) we obtain
$\mathrm{KL}(\mu_{n\gamma}||\pi)$ $=\frac{1}{2}\left(-\log\det(I+A_{n})+Tr[A_{n}]+Tr[A_{0}^{-1}C_{\pi}^{-1}\varepsilon_{0}\varepsilon_{0}^{\top}A_{0}^{-1}e^{n\gamma C_{\pi}^{-1}}A_{n}^{2}e^{n\gamma C_{\pi}^{-1}}]\right)$
$=\frac{1}{2}\left(-\log\det(I+A_{n})+Tr[A_{n}]+Tr[F_{n}A_{n}^{2}]\right)$
where $F_{n}:=e^{n\gamma C_{\pi}^{-1}}D_{0}e^{n\gamma C_{\pi}^{-1}}$ and $D_{0}:=A_{0}^{-1}C_{\pi}^{-1}\varepsilon_{0}\varepsilon_{0}^{\top}A_{0}^{-1}$ and recall that using (3.10)
$A_{n}$ $=e^{-n\gamma\Gamma}\left(A_{0}^{-1}+\Omega\,(I-e^{-2n\gamma\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$
Denote $K_{n}:=\left(A_{0}^{-1}+\Omega\,(I-e^{-2n\gamma\Gamma})\right)^{-1}$ so that $A_{n}=e^{-n\gamma\Gamma}K_{n}e^{-n\gamma\Gamma}$. Since $A_{n}=C_{\pi}^{-1}C_{n\gamma}-I$ has eigenvalues larger than 1-1, we have $\det(I+A_{n})=e^{Tr[\log(I+A_{n})]}$, and
$-\log\det(I+A_{n})=-Tr[\log(I+A_{n})]=\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}Tr[A_{n}^{k}]$
whenever $A_{n}||<1$ (so that the series converges), which holds for sufficiently large $n$. Then
$2\mathrm{KL}(\mu_{n\gamma}||\pi)=Tr[F_{n}A_{n}^{2}]+\frac{1}{2}Tr[A_{n}^{2}]+\sum_{k=3}^{\infty}\frac{(-1)^{k}}{k}Tr[A_{n}^{k}]$
where $\sum_{k=3}^{\infty}\frac{(-1)^{k}}{k}Tr[A_{n}^{k}]$ is convergent and equal to $\log\det(I+A_{n})-Tr[A_{n}^{2}]/2+Tr[A_{n}]$. Also, $\Gamma$ is diagonalisable when $C_{\pi}$ is positive definite and we denote its spectral decomposition by $\Gamma=P^{\top}\Lambda P$ with $\lambda_{i}=\Lambda_{ii}$ denoting the $i$th eigenvalue of $\Gamma$ and $\frac{1}{2}<\lambda_{1}<\lambda_{2}<\lambda_{3}\dots<\lambda_{d}$ and $p_{i}$ the $i$th column of $P$ where $P$ is unitary. Then
$Tr[A_{n}^{2}]$ $=Tr[(P^{\top}e^{-n\gamma\Lambda}PK_{n}P^{\top}e^{-n\gamma\Lambda}P)^{2}]$
$=Tr[e^{-2n\gamma\Lambda}Me^{-2n\gamma\Lambda}M]$
$=\sum_{i=1}^{d}\sum_{j=1}^{d}e^{-2n\gamma(\lambda_{i}+\lambda_{j})}M_{ij}M_{ji}$
where $M:=PK_{n}P^{\top}$ since $Tr[P^{\top}DP]=\sum_{i=1}^{d}D_{ii}$ for $P$ unitary orthonormal matrix, $PP^{\top}=I$. A similar representation can be obtained for $Tr[A_{n}^{k}]$ involving $k$ iterated sums. Similarly,
$Tr[F_{n}A_{n}^{2}]$ $=e^{-n\gamma}Tr[K_{n}D_{0}K_{n}e^{-2n\gamma\Gamma}]=e^{-n\gamma}Tr[e^{-2n\gamma\Lambda}PK_{n}D_{0}K_{n}P^{\top}]=\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+\frac{1}{2})}N_{ii}$
where $N:=PK_{n}D_{0}K_{n}P^{\top}$. Using the same reasoning for $\mathrm{KL}(\nu_{n}||\pi)$ yields
$\frac{\mathrm{KL}(\nu_{n}||\pi)}{\mathrm{KL}(\mu_{n\gamma}||\pi)}$ $=\frac{\frac{1}{2}\sum_{i=1}^{d}\sum_{j=1}^{d}e^{-2n\gamma(\lambda_{i}+\lambda_{j})}M_{ij}^{\beta}M_{ji}^{\beta}+\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+\frac{1}{2})}N_{ii}^{\beta}+\sum_{k=3}^{\infty}\frac{(-1)^{k}}{k}Tr[B_{n}^{k}]}{\frac{1}{2}\sum_{i=1}^{d}\sum_{j=1}^{d}e^{-2n\gamma(\lambda_{i}+\lambda_{j})}M_{ij}M_{ji}+\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+\frac{1}{2})}N_{ii}+\sum_{k=3}^{\infty}\frac{(-1)^{k}}{k}Tr[A_{n}^{k}]}$
$=\frac{\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+\frac{1}{2})}N_{ii}^{\beta}+\mathcal{O}(e^{-4n\gamma\lambda_{1}})}{\sum_{i=1}^{d}e^{-2n\gamma(\lambda_{i}+\frac{1}{2})}N_{ii}+\mathcal{O}(e^{-4n\gamma\lambda_{1}})}$
$=\frac{N_{11}^{\beta}+\mathcal{O}(e^{-n\gamma\min(4\lambda_{1},2(\lambda_{2}-\lambda_{1}))})}{N_{11}+\mathcal{O}(e^{-n\gamma\min(4\lambda_{1},2(\lambda_{2}-\lambda_{1}))})}$
where $M^{\beta}=PK_{n}^{\beta}P^{\top}$, $N^{\beta}:=PK_{n}^{\beta}D_{0}K_{n}^{\beta}P^{\top}$, $K_{n}^{\beta}:=(A_{0}^{-1}+\Omega^{\beta}(I-e^{-2n\gamma\Gamma}))^{-1}$ and $B_{n}=e^{-n\gamma\Gamma}K_{n}^{\beta}e^{-n\gamma\Gamma}$ and in the third line we have divided top and bottom by $e^{-2n\gamma(\lambda_{1}+\frac{1}{2})}$ and noting that $K_{n}\rightarrow(A_{0}^{-1}+\Omega)^{-1}$ and $K_{n}^{\beta}\rightarrow(A_{0}^{-1}+\Omega^{\beta})^{-1}$ as $n\rightarrow\infty$. Taking the limit of the above expression as $n\rightarrow\infty$ yields the result.
Appendix C Preservation of log-concavity
Our proof strategy involves starting with a sequential splitting of the WFR flow and using Girsanov theorem to characterise the intermediate density due to the W flow. This is based on a similar proof strategy used to show that the filtering density satisfies a Poincaré inequality (see Lemma 5.1 in [Pathiraja2021]). In Appendix C.1, we show that when the initial density is “sufficiently” strongly log-concave (to be made precise), there exists a finite time interval over which the W flow maintains log-concavity. Although the W flow cannot be expected to preserve log-concavity uniformly in time (unless $\pi$ is Gaussian), the properties of the FR flow can be exploited to maintain log-concavity uniformly in time. The next lemma shows that the FR does exactly this, and then the proof in Appendix C.2 combines the results.
Lemma C.1.
FR flows preserve log-concavity. Suppose $\mu_{t}$ is the solution of the FR flow (2.3) at time $t$, initialised at $\mu_{0}(x)$ satisfying Assumption 2. Then $\mu_{t},\enskip t>0$ is $\alpha_{t}$-strongly log-concave with
$\alpha_{t}=(1-e^{-t})\alpha_{\pi}+e^{-t}\alpha_{0}$
Proof.
Using the exact solution to the FR flow (2.5),
$\mu_{t}(x)\propto\pi(x)^{1-e^{-t}}\mu_{0}(x)^{e^{-t}}\propto e^{-V_{t}(x)}$
where $V_{t}=(1-e^{-t})V_{\pi}+e^{-t}\mu_{0}$. Also
$\nabla^{2}V_{t}=(1-e^{-t})\nabla^{2}V_{\pi}+e^{-t}\nabla^{2}V_{0}\succeq((1-e^{-t})\alpha_{\pi}+e^{-t}\alpha_{0})I=:\alpha_{t}I$
where $\alpha_{t}>0$ is a convex combination of positive scalars for all $t>0$. ∎
C.1 Proof of Lemma 4.1
Proof.
The first step will be to use Girsanov theorem to characterise the law of the overdamped Langevin diffusion, using the re-scaled Langevin diffusion
$dX_{t}=\nabla U(X_{t})dt+dW_{t}$
where $U=-\frac{1}{2}V_{\pi}$. Let $\widehat{\mu}_{t}$ denote the law of $X_{t}$, so then
$\widehat{\mu}_{t}$ $=\int_{C[0,t]}\exp\left(\int_{0}^{t}\nabla U(W_{s})dWs-\frac{1}{2}\int_{0}^{t}|\nabla U(W_{s})|^{2}ds\right)\mu_{W}(dw)$
$=\int_{C[0,t]}\exp\left((U(W_{t})-U(W_{0})-\frac{1}{2}\int_{0}^{t}(\Delta U+|\nabla U|^{2})(W_{s})ds\right)\mu_{W}(dw)$
where $\mu_{W}(dw)$ denotes the Wiener measure on path space $C[0,t]$ and in the second line we have used Itô formula. Here the application of Girsanov theorem is valid since Novikov’s condition is satisfied under Assumption 1. Consider now a discretisation of the time interval $[0,t]$ with step size $\tau$ and $N$ steps, $0=t_{0},t_{1},t_{2},\dots,t_{N-1},t_{N}=t$ with $t_{i}-t_{i-1}=\tau$ for all $i=1,2,3,\dots,N$. From now on, we use the notation $\widehat{\mu}_{i}$ to denote an approximation to $\widehat{\mu}_{t_{i}}$ (and likewise for other quantities). We can then approximate as
$\widehat{\mu}_{t}\approx$ $\int_{C[0,t]}\exp\left(\frac{1}{2}(V_{\pi}(w_{0})-V_{\pi}(w_{N}))\right)\prod_{i=1}^{N-1}\exp\left(-\frac{1}{2}R(w_{i})\right)\mu_{W}(dw)$
$=\exp\left(-\frac{1}{2}V_{\pi}(w_{N})\right)\prod_{i=1}^{N-1}\mu_{i}(w_{i})$
where
$\mu_{i}(w_{i}):=\left\{\begin{array}[]{lr}\int Q(w_{{i-1}})\,q_{\tau}(w_{{i-1}},w_{i})\,\mu_{i-1}(w_{i-1})\,dw_{{i-1}},&i=2,3,\dots\\ \int Q(w_{0})\,q_{\tau}(w_{0},w_{i})\,\exp\left(-U(w_{0})\right)\,\exp(-V_{0}(w_{0}))\,dw_{0},&i=1\end{array}\right$.
and
$Q(w)$ $:=\exp\left(-\frac{1}{2}R(w)\tau\right)$
$R(w)$ $:=\Delta U(w)+|\nabla U(w)|^{2}$
$q_{\tau}(w_{{i-1}},w_{i})$ $:=\frac{1}{(2\pi\tau)^{d/2}}\exp\left(-\frac{|w_{i}-w_{{i-1}}|^{2}}{2\tau}\right)$
Then by Assumption 3, for $i=1$, we have,
$\mu_{1}$ $\propto\int\exp\left(-\frac{|w_{0}-w_{1}|^{2}}{2\tau}+\frac{1}{2}V_{\pi}(w_{0})-\frac{\tau}{2}R(w_{0})-V_{0}(w_{0})\right)dw_{0}$
$=\int\exp\left(-\frac{|w_{0}-w_{1}|^{2}}{2\tau}-\left(V_{0}(w_{0})-\frac{1+\delta}{2}V_{\pi}(w_{0})\right)-\frac{\delta-\tau}{2}V_{\pi}(w_{0})-\frac{\tau}{2}\left(R(w_{0})+V_{\pi}(w_{0})\right)\right)dw_{0}$
$=:\int\exp(-f_{1}(w_{1},w_{0}))dw_{0}$
$=\exp(-\mathcal{G}_{1}(w_{1}))$
assuming $\tau<\delta$. Since $f_{1}$ is strongly convex in both $w_{1}$ and $w_{0}$ under Assumption 3, by the Prékopa–Leindler inequality [saumard2014log, Theorem 3.8], so is $\mathcal{G}_{1}$ with
$\nabla_{y}^{2}\mathcal{G}_{1}\succeq\frac{\int(\nabla_{y}^{2}f_{1}-\nabla_{yz}^{2}f_{1}(\nabla_{z}^{2}f_{1})^{-1}\nabla_{zy}^{2}f_{1})\exp(-f_{1}(y,z))dz}{\int\exp(-f_{1}(y,z))dz},$ (C.1)
using the shorthand notation $y:=w_{1}$ and $z:=w_{0}$. Then,
$\nabla_{w_{1}}^{2}f_{1}$ $=\frac{1}{\tau}I$
$\nabla_{w_{1}w_{0}}^{2}f_{1}$ $=\nabla_{w_{0}w_{1}}^{2}f_{1}=-\frac{1}{\tau}I$
$\nabla_{w_{0}}^{2}f_{1}$ $=\nabla_{w_{0}}^{2}(V_{0}-\frac{1+\delta}{2}V_{\pi})+\frac{\tau}{2}\nabla_{w_{0}}^{2}\mathcal{H}+\frac{\delta-\tau}{2}\nabla_{w_{0}}^{2}V_{\pi}+\frac{1}{\tau}I$
$\succeq\left(\alpha_{d}+\frac{\tau}{2}\alpha_{h}+\frac{\delta-\tau}{2}\alpha_{\pi}+\frac{1}{\tau}\right)I$
So then
$\nabla_{y}^{2}\mathcal{G}_{1}$ $\succeq\frac{\int(\frac{1}{\tau}-\frac{1}{\tau^{2}}(\nabla_{w_{0}}^{2}f)^{-1})\exp(-f(y,z))dz}{\int\exp(-f(y,z))dz}$
$\succeq\left(\frac{1}{\tau}-\frac{1}{\tau^{2}}\left(\alpha_{d}+\frac{\tau}{2}\alpha_{h}+\frac{\delta-\tau}{2}\alpha_{\pi}+\frac{1}{\tau}\right)^{-1}\right)I\frac{\int\exp(-f(y,z))dz}{\int\exp(-f(y,z))dz}$
$=\frac{\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}+\frac{\tau}{2}(\alpha_{h}-\alpha_{\pi})}{1+\tau(\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}+\frac{\tau}{2}(\alpha_{h}-\alpha_{\pi}))}I$ (C.2)
$=:c_{1}I$
which is strictly positive by all the assumptions and since $\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}+\frac{\tau}{2}(\alpha_{h}-\alpha_{\pi})=\alpha_{d}+\frac{\tau}{2}\alpha_{h}+\frac{\delta-\tau}{2}\alpha_{\pi}>0$ whenever $\delta>\tau$ . Then for $i=2$,
$\mu_{2}$ $\propto\int\exp\left(-\frac{|w_{1}-w_{2}|^{2}}{2\tau}-\frac{\tau}{2}R(w_{1})-\mathcal{G}_{1}(w_{1})\right)dw_{1}$
$=\int\exp\left(-\frac{|w_{1}-w_{2}|^{2}}{2\tau}-\frac{\tau}{2}(R(w_{1})+V_{\pi}(w_{1}))+\frac{\tau}{2}V_{\pi}(w_{1})-\mathcal{G}_{1}(w_{1})\right)dw_{1}$
$=:\int\exp\left(-f_{2}(w_{1},w_{2})\right)dw_{1}.$
Once again, $f_{2}$ is strongly convex in $w_{1}$ and $w_{2}$ and also
$\nabla_{w_{1}}^{2}f_{2}$ $=\frac{1}{\tau}I+\frac{\tau}{2}\nabla_{w_{1}}^{2}\mathcal{H}(w_{1})-\frac{\tau}{2}\nabla_{w_{1}}^{2}V_{\pi}(w_{1})+\nabla_{w_{1}}^{2}\mathcal{G}_{1}(w_{1})$
$\succeq\left(\frac{1}{\tau}+\frac{\tau}{2}\alpha_{h}+c_{1}-\frac{\tau}{2}L_{\pi}\right)I.$
A sufficient condition to maintain convexity is to choose $\tau$ small enough such that $c_{1}-\frac{\tau}{2}L_{\pi}>0$ (which is possible since $c_{1}\rightarrow\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}>0$ as $\tau\rightarrow 0$). Then again by Prékopa–Leindler, $\mu_{2}\propto\int\exp(-f_{2}(w_{1},w_{2}))dw_{1}=\exp(-\mathcal{G}_{2}(w_{2}))$ with
$\nabla^{2}_{w_{2}}\mathcal{G}_{2}$ $\succeq\left(\frac{1}{\tau}-\frac{1}{\tau^{2}}\left(\frac{1}{\tau}+\frac{\tau}{2}\alpha_{h}+\kappa_{1}(\tau)\right)^{-1}\right)I$
$=\frac{\kappa_{1}(\tau)+\frac{\tau}{2}\alpha_{h}}{1+\tau(\kappa_{1}(\tau)+\frac{\tau}{2}\alpha_{h})}I=:c_{2}I$
where $\kappa_{1}(\tau):=c_{1}-\frac{\tau}{2}L_{\pi}$. Repeating for $\mu_{3}$, we have exactly the same computations, but with the requirement that $\tau$ is chosen small enough that $c_{2}-\frac{\tau}{2}L_{\pi}>0$, which by similar arguments as previously, holds true for some sufficiently small $\tau$. By induction, we conclude that for all $i=1,2,3,\dots,N$, $\mu_{i}\propto\exp(-\mathcal{G}_{i}(w_{i}))$ is strongly log-concave with $\nabla_{w_{i}}^{2}\mathcal{G}_{i}\succeq c_{i}I$ where
$c_{i}$ $=\frac{c_{i-1}+\frac{\tau}{2}(\alpha_{h}-L_{\pi})}{1+\tau(c_{i-1}+\frac{\tau}{2}(\alpha_{h}-L_{\pi}))},\quad i=1,2,3,\dots,N$ (C.3)
$c_{0}$ $=\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}.$ (C.4)
Finally, returning to our original approximation, we have that
$\widehat{\mu}_{t}\approx$ $\exp\left(-\frac{1}{2}V_{\pi}(w_{N})\right)\mu_{N}(w_{N})\propto\exp\left(-\frac{1}{2}V_{\pi}(w_{N})-\mathcal{G}_{N}\right)$
and
$\frac{1}{2}V_{\pi}(w_{N})+\mathcal{G}_{N}(w_{N})\succeq\left(\frac{1}{2}\alpha_{\pi}+c_{N}\right)I.$ (C.5)
In order to understand the limit $\tau\rightarrow 0$ of the recursion (C.3), notice that it can be seen as a two step time discretisation, i.e.
$\tilde{c}_{i}$ $=c_{i-1}-\tau b^{2}$ (C.6)
$c_{i}$ $=\frac{\tilde{c}_{i}}{1+\tau\tilde{c}_{i}}$ (C.7)
with $b^{2}=\frac{|\alpha_{h}-L_{\pi}|}{2}$, recalling that $\alpha_{h}-L_{\pi}<0$. Also, (C.6) corresponds to an Euler discretisation of the ODE $\frac{dc_{t}}{dt}=-b^{2}$ with time step $\tau$. Notice also that (C.7) is the exact solution of $\frac{dc_{t}}{dt}=-c_{t}^{2}$ initialised at $\widehat{c}_{i}$ at time $t=\tau$ since if $c_{t}=c_{0}(1+tc_{0})^{-1}$ then differentiating with respect to $t$ yields $\dot{c}_{t}=-c_{0}^{2}(1+tc_{0})^{-2}=-c_{t}^{2}$. Therefore, the iteration (C.6)-(C.7) corresponds to an operator splitting based discretisation of the ODE
$\frac{dc_{t}}{dt}=-c_{t}^{2}-b^{2}.$ (C.8)
Notice that for the case $\frac{dc_{t}}{dt}=-c_{t}^{2}-b^{2}$, the ODE has no stable fixed point, and continues decreasing to -$\infty$ as $t\rightarrow\infty$, consistent with the findings of [Kolesnikov2001] that the W flow does not preserve strong log-concavity uniformly in time (aside from the Ornstein-Ulenbeck flow). We will then focus on characterising the time horizon over which strong convexity is preserved. Integrating both sides of the ODE and assuming $c_{t}^{2}+b^{2}\neq 0$ for all $t$ yields
$c_{t^{\ast}}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-bt^{\ast}\right).$
To summarise, the exact solution to (C.8) is given by
$c_{t}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-bt\right),$ (C.9)
$c_{0}$ $=\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}$ (C.10)
where $b:=\sqrt{\frac{|\alpha_{h}-L_{\pi}|}{2}}$, noting this excludes the final $\frac{1}{2}\alpha_{\pi}$ term. Finally, we have that
$\widehat{\mu}_{t}$ $\propto\exp\left(-\frac{1}{2}V_{\pi}-\mathcal{G}_{t}\right)$
$=:\exp\left(-\mathcal{E}_{t}\right)$
$\nabla^{2}\mathcal{E}_{t}$ $\succeq\left(c_{t}(c_{0})+\frac{1}{2}\alpha_{\pi}\right)I.$
When $\alpha_{h}-L_{\pi}<0$, there exists some $t^{\ast}$ beyond which strong convexity111of the intermediate densities is lost, that is, $c_{t}\leq 0$ for all $t\geq t^{\ast}$. Due to the monotonocity of the tan function, $t^{\ast}$ is easily obtained as
$t^{\ast}$ $=\frac{1}{b}\tan^{-1}\left(\frac{c_{0}}{b}\right)-\frac{1}{b}\tan^{-1}\left(0\right)$ (C.11)
$=\frac{1}{b}\tan^{-1}\left(\frac{c_{0}}{b}\right)$ (C.12)
which is strictly positive since $b,c_{0}>0$.
C.2 Proof of Theorem 4.1
Proof.
Consider a fixed time interval $[0,T]$. We will characterise strong log-concavity for $t\in[0,T]$ under (2.1) and then demonstrate that this holds for all $T>0$.
Consider the sequence of densities due to the W-FR splitting scheme (2.6). Recall that $\widehat{\nu}_{i}$ is the density after a step of W flow and $\mu_{i}$ is the density after a step of FR flow. We first show that strong log-concavity is preserved for a sequence of steps with step size $\{\tau_{i}\}_{i=1,2,\dots}$, and then show that strong log-concavity is preserved as $\tau_{i}\rightarrow 0$. As will become clear in the remainder of the proof, such an iteration dependent step size is needed due to the fact that the time horizon over which the W flow preserves strong log-concavity depends on the relative convexity of the initial and target potentials.
For a given strongly log-concave $\mu_{0}$ satisfying Assumptions 3, we have that after a single step of the W flow of size $\tau_{1}$, the distribution $\widehat{\nu}_{1}$ is $\widehat{\alpha}_{1}$-strongly log-concave due to Lemma 4.1 with potential $\mathcal{E}_{1}=\frac{1}{2}V_{\pi}+\mathcal{G}_{1}$, where $\nabla^{2}\mathcal{E}_{1}\succeq\widehat{\alpha}_{1}I$ and $\nabla^{2}\mathcal{G}_{1}\succeq\widehat{c}_{1}I$, $\widehat{\alpha}_{1}=\widehat{c}_{1}+\frac{\alpha_{\pi}}{2}$ and
$\widehat{c}_{1}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-b\tau_{1}\right)$
$c_{0}$ $=\alpha_{d}+\frac{\delta_{0}}{2}\alpha_{\pi},$
where $0<\delta_{0}<1$ and $\tau_{1}$ must be “small enough” that $\widehat{c}_{1}>0$, and due to Lemma 4.1 and the assumed conditions, such a $\tau_{1}>0$ exists. For the FR step, it trivially holds that
$\mu_{1}$ $\propto\pi^{1-e^{-\tau_{1}}}\widehat{\mu}_{1}^{e^{-\tau_{1}}}\propto\exp\left(-(1-e^{-\tau_{1}})V_{\pi}-e^{-\tau_{1}}\mathcal{E}_{1}\right)=:\exp(-\mathcal{F}_{1})$
where $\mathcal{F}_{1}=V_{\pi}+e^{-\tau_{1}}(\mathcal{E}_{1}-V_{\pi})=V_{\pi}+e^{-\tau_{1}}(\mathcal{G}_{1}-\frac{1}{2}V_{\pi})$ yields that the distribution from a single sequential split step of size $\tau_{1}$ is strongly log-concave with constant
$\alpha_{1}=\alpha_{\pi}+e^{-\tau_{1}}(\widehat{\alpha}_{1}-\alpha_{\pi})=\left(1-\frac{1}{2}e^{-\tau_{1}}\right)\alpha_{\pi}+e^{-\tau_{1}}\widehat{c}_{1}.$
For $i=2$, we require that $\mathcal{F}_{1}-\frac{1+\delta_{1}}{2}V_{\pi}$ is strongly convex to apply Lemma 4.1, for some $0<\delta_{1}<1$. Note that we use a different $\delta_{1}$ as compared to $\delta_{0}$ in the first step, as this will be used to define the sequence of step sizes. Then
$\nabla^{2}(\mathcal{F}_{1}-\tfrac{1+\delta_{1}}{2}V_{\pi})$ $=\nabla^{2}\left(\left(1-\tfrac{1+\delta_{1}}{2}\right)V_{\pi}+e^{-\tau_{1}}(\mathcal{G}_{1}-\tfrac{1}{2}V_{\pi})\right),$
$=\nabla^{2}\left(\left(\tfrac{1-\delta_{1}-e^{-\tau_{1}}}{2}\right)V_{\pi}+e^{-\tau_{1}}\mathcal{G}_{1}\right)$
$\succeq\left(\left(\tfrac{1-\delta_{1}-e^{-\tau_{1}}}{2}\right)\alpha_{\pi}+e^{-\tau_{1}}\widehat{c}_{1}\right)I$
$=:\alpha_{d,1}I$
where $\alpha_{d,1}$ denotes the concavity parameter required for Assumption 3(a) in Lemma 4.1. A sufficient condition for strict positivity of $\alpha_{d,1}$ is to ensure $0<\delta_{1}<1-e^{-\tau_{1}}$. Then $\widehat{\nu}_{2}$ is again strongly convex by Lemma 4.1 with potential $\mathcal{E}_{2}=\frac{1}{2}V_{\pi}+\mathcal{G}_{2}$, $\nabla^{2}\mathcal{G}_{2}\succeq\widehat{c}_{2}I$ and
$\widehat{c}_{2}$ $=b\tan\left(\tan^{-1}\left(\tfrac{c_{1}}{b}\right)-b\tau_{2}\right)$
$c_{1}$ $=\left(\tfrac{1-\delta_{1}-e^{-\tau_{1}}}{2}\right)\alpha_{\pi}+e^{-\tau_{1}}\widehat{c}_{1}+\tfrac{\delta_{1}}{2}\alpha_{\pi}$
$=\left(\tfrac{1-e^{-\tau_{1}}}{2}\right)\alpha_{\pi}+e^{-\tau_{1}}\widehat{c}_{1}$
$=\tfrac{1}{2}\alpha_{\pi}+e^{-\tau_{1}}(\widehat{c}_{1}-\tfrac{\alpha_{\pi}}{2})$
and $c_{1}>0$ since $\tau_{1}$ is chosen such that $\widehat{c}_{1}>0$. Once again, $\tau_{2}$ must be chosen such that $\widehat{c}_{2}>0$. Then once again for the FR step, we have $\nu_{2}\propto\exp(-\mathcal{F}_{2})$ with
$\mathcal{F}_{2}$ $=(1-e^{-\tau_{2}})V_{\pi}+e^{-\tau_{2}}\mathcal{E}_{2}$
$=V_{\pi}+e^{-\tau_{2}}(\mathcal{E}_{2}-V_{\pi})$
$=V_{\pi}+e^{-\tau_{2}}(\mathcal{G}_{2}-\frac{1}{2}V_{\pi})$
and $\nabla^{2}\mathcal{F}_{2}\succ\alpha_{2}I$ where
$\alpha_{2}=\left(1-\frac{1}{2}e^{-\tau_{2}}\right)\alpha_{\pi}+e^{-\tau_{2}}\widehat{c}_{2}.$
Proceeding inductively for $i=3,4,\dots$, in a similar way with required conditions on $\tau_{i}$ and $\delta_{i}$, we obtain
$\alpha_{i}$ $=\left(1-\tfrac{1}{2}e^{-\tau_{i}}\right)\alpha_{\pi}+e^{-\tau_{i}}\widehat{c}_{i}=c_{i}+\frac{\alpha_{\pi}}{2},\quad i=1,2,3,\dots$ (C.13)
where $\widehat{c}_{{i}}$ and $c_{i}$ satisfy the recursion for $i=1,2,3,\dots$
$\widehat{c}_{i}$ $=b\tan\left(\tan^{-1}\left(\tfrac{c_{i-1}}{b}\right)-b\tau_{i}\right)$ (C.14)
$c_{i}$ $=(1-e^{-\tau_{i}})\frac{\alpha_{\pi}}{2}+e^{-\tau_{i}}\widehat{c}_{i}$ (C.15)
$c_{0}$ $=\alpha_{d}+\frac{\delta_{0}}{2}\alpha_{\pi}.$ (C.16)
Similarly to the limiting analysis in the proof of Lemma 4.1, we have that (C.14)-(C.15) corresponds to a splitting scheme of the ODE
$\dot{c}_{t}=-c_{t}^{2}-c_{t}-b^{2}+\frac{\alpha_{\pi}}{2},$ (C.17)
whereby for the $i$th iteration, the ODE $\dot{c}_{t}=-c_{t}^{2}-b^{2}$, initialised at $c_{{i-1}}$ is solved over time $\tau_{i}$ yielding $\widehat{c}_{i}$ and then the ODE $\dot{c}_{t}=-c_{t}+\alpha_{\pi}$ initialised at $\widehat{c}_{i}$ is solved over time $\tau_{i}$ yielding $c_{i}$. Finally, due to (C.13), $\alpha_{t}=c_{t}+\frac{\alpha_{\pi}}{2}$, is the log-concavity constant of $\mu_{t}$, the solution of the exact WFR PDE (2.1) at time $t$.
We now make precise the conditions on $\tau_{i}$ which must be satisfied in order for the limit (C.17) to hold as $\max_{i}\tau_{i}\rightarrow 0$. Recall that $\tau_{i}$ must be chosen such that $\widehat{c}_{i}>0$, which using (C.12), is possible if $0<\tau_{i}<\tau_{i}^{\ast}$ where
$\tau_{i}^{\ast}$ $=\frac{1}{b}\tan^{-1}\left(\frac{c_{i-1}}{b}\right),\quad i=1,2,3,\dots$ (C.18)
Notice that there exists a $\tau_{1}^{\ast}>0$ whenever $\alpha_{d},\delta_{0}>0$, which implies $\widehat{c}_{1}>0$ and $c_{1}>\left(1-e^{-\tau_{1}}\right)\frac{\alpha_{\pi}}{2}$ and $\tau_{2}^{\ast}>\frac{1}{b}\tan^{-1}\left(\frac{(1-e^{-\tau_{1}})\alpha_{\pi}}{2b}\right)>0$. This then implies $\widehat{c}_{2}>0$ and once again $c_{2}>\left(1-e^{-\tau_{2}}\right)\frac{\alpha_{\pi}}{2}$ and $\tau_{3}^{\ast}>\frac{1}{b}\tan^{-1}\left(\frac{(1-e^{-\tau_{2}})\alpha_{\pi}}{2b}\right)>0$. By induction, $\tau_{i}^{\ast}>\frac{1}{b}\tan^{-1}\left(\frac{(1-e^{-\tau_{i-1}})\alpha_{\pi}}{2b}\right)>0$ and since $\tau_{i-1}<\tau_{i-1}^{\ast}$, we can instead consider the recursion
$\tau^{\ast}_{i}=\frac{1}{b}\tan^{-1}\left(\frac{(1-e^{-\tau_{i-1}^{\ast}})\alpha_{\pi}}{2b}\right),\enskip i=1,2,3,\dots$ (C.19)
It is not difficult to see that $\tau_{i}^{\ast}>0$ for all $i=1,2,3,\dots$. Its limit, $\tau_{\infty}^{\ast}$ can be found from the fixed point equation $\tau^{\ast}_{\infty}=g(\tau_{\infty}^{\ast})$ where $g(\tau):=\frac{1}{b}\tan^{-1}\left(\frac{(1-e^{-\tau})\alpha_{\pi}}{2b}\right)$, for which $t_{\infty}^{\ast}=0$ is a valid (but not unique) solution. Notice also that $0<\frac{dg(\tau)}{d\tau}=\frac{\alpha_{\pi}}{2b^{2}}\frac{e^{-\tau}}{1+\left(\frac{(1-e^{-\tau})\alpha_{\pi}}{2b}\right)^{2}}<\frac{\alpha_{\pi}}{2b^{2}}$ for all $\tau>0$. Under the assumed condition (4.4), $\frac{dg(0)}{d\tau}=\frac{\alpha_{\pi}}{2b^{2}}>0$, $\tau=0$ is an unstable fixed point, so that $\tau_{\infty}^{\ast}\neq 0$ since $\tau_{1}^{\ast}\neq 0$. To show that there exists at least one other $0<\tau_{\infty}^{\ast}<\infty$, consider $h(\tau):=g(\tau)-\tau$. First note that $h(\tau)\rightarrow$-$\infty$ as $\tau\rightarrow\infty$ since $g(\tau)<\frac{1}{b}\tan^{-1}\left(\frac{\alpha_{\pi}}{2b}\right)$ for all $\tau\geq 0$. Additionally, $h(0)=0$ and $\frac{dh(0)}{d\tau}>0$ and $\frac{dh(\tau)}{d\tau}\rightarrow-1$ as $\tau\rightarrow\infty$ monotonically.
Then by continuity of $h$ and $g$, there must exist a $\tau>0$ such that $h(\tau)=0$, i.e. there exists a $\tau_{\infty}^{\ast}>0$. Therefore, we may construct a sequence of step sizes $\{\tau_{i}\}_{i=1,2,3,}$ with $\tau_{i}<\min(\tau_{1}^{\ast},\tau_{\infty}^{\ast}\neq 0)$ for all $i=1,2,3,\dots$ as a discretisation of the interval $[0,T]$ for all $T>0$, for which the recursion (C.14)-(C.15) converges to the continuous time ODE (C.17) as $\max_{i}\tau_{i}\rightarrow 0$.
As a further confirmation of the need for condition (4.4), consider the case $\frac{\alpha_{\pi}}{2b^{2}}<1$. The recursion (C.19) is a decreasing sequence with $\tau_{\infty}^{\ast}=0$ by a standard fixed point argument. Furthermore, the bound on $\frac{dg(\tau)}{d\tau}$ implies $g$ is Lipschitz and so $|\tau_{i+1}^{\ast}-\tau_{i}^{\ast}|=|g(\tau_{i}^{\ast})-g(\tau_{i-1}^{\ast})|\leq\frac{\alpha_{\pi}}{2b^{2}}|\tau_{i}^{\ast}-\tau_{i-1}^{\ast}|$. Then since $g(0)=0$, it holds that $\tau_{i}^{\ast}\leq\left(\frac{\alpha_{\pi}}{2b^{2}}\right)^{i}\tau_{0}^{\ast}$. Since $\sum_{i=1}^{\infty}\left(\frac{\alpha_{\pi}}{2b^{2}}\right)^{i}=\frac{1}{1-\frac{\alpha_{\pi}}{2b^{2}}}$ whenever $\frac{\alpha_{\pi}}{2b^{2}}<1$, it holds that $\sum_{i=1}^{\infty}\tau_{i}<\sum_{i=1}^{\infty}\tau_{i}^{\ast}\leq\frac{1}{1-\frac{\alpha_{\pi}}{2b^{2}}}$, so that a uniform in time control does not follow from the above sequence of thresholds $\{\tau_{i}^{\ast}\}$.
Finally, to obtain a uniform in time log-concavity constant, consider the fixed point of (C.17),
$c_{\infty}=\frac{-1\pm\sqrt{1+4(\frac{\alpha_{\pi}}{2}-b^{2})}}{2},$ (C.20)
which has one real positive solution under condition (4.4). Note also that $c_{t}$ approaches $c_{\infty}$ monotonically, as is true for scalar, autonomous Riccati ODEs. Its explicit solution is trivially found by solving the separable ODE, which we include here for completeness. Denote by $c_{\infty,1}$ and $c_{\infty,2}$, the positive and negative solutions of (C.20), respectively. Then,
$\int_{c_{0}}^{c_{T}}\frac{1}{(c_{t}-c_{\infty,1})(c_{t}-c_{\infty,2})}dc_{t}$ $=-\int_{0}^{T}dt$
$\frac{(c_{T}-c_{\infty,1})(c_{0}-c_{\infty,2})}{(c_{T}-c_{\infty,2})(c_{0}-c_{\infty,1})}$ $=e^{-T(c_{\infty,1}-c_{\infty,2})}$
$(c_{T}-c_{\infty,1})$ $=l_{0}e^{-T(c_{\infty,1}-c_{\infty,2})}(c_{T}-c_{\infty,2})$
$c_{T}$ $=\frac{c_{\infty,1}-c_{\infty,2}l_{0}e^{-T(c_{\infty,1}-c_{\infty,2})}}{1-l_{0}e^{-T(c_{\infty,1}-c_{\infty,2})}}$
where $l_{0}=\frac{(c_{0}-c_{\infty,1})}{(c_{0}-c_{\infty,2})}$. Finally, noting that $\alpha_{T}=c_{T}+\frac{\alpha_{\pi}}{2}$ yields the result.